### v3

# Enhancing High-Level Synthesis with Automated Pragma Insertion and Code Transformation Framework

Abstract.
High-level synthesis, source-to-source compilers, and various Design Space Exploration techniques for pragma insertion have significantly improved the Quality of Results of generated designs. These tools offer benefits such as reduced development time and enhanced performance. However, achieving high-quality results often requires additional manual code transformations and tiling selections, which are typically performed separately or as pre-processing steps. Although DSE techniques enable code transformation upfront, the vastness of the search space often limits the exploration of all possible code transformations, making it challenging to determine which transformations are necessary. Additionally, ensuring correctness remains challenging, especially for complex transformations and optimizations.
To tackle this obstacle, we first propose a comprehensive framework leveraging HLS compilers. Our system streamlines code transformation, pragma insertion, and tiles size selection for on-chip data caching through a unified optimization problem, aiming to enhance parallelization, particularly beneficial for computation-bound kernels. Them employing a novel Non-Linear Programming (NLP) approach, we simultaneously ascertain transformations, pragmas, and tile sizes, focusing on regular loop-based kernels. Our evaluation demonstrates that our framework adeptly identifies the appropriate transformations, including scenarios where no transformation is necessary, and inserts pragmas to achieve a favorable Quality of Results.
HLS, code transformation, pragma insertion, non-linear problem
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; October 27–31, 2024; New Jersey, USAisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
High-level synthesis (HLS) compilers and source-to-source compilers are indispensable tools in accelerating hardware design by automating the translation of high-level programming languages like C/C++ or Python into hardware descriptions. They offer notable advantages, such as reduced development time and enhanced performance for hardware designs (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020). However, achieving a good quality of results (QoR) often requires manual code transformations and pragma insertions, which can be facilitated with either Design Space Exploration (DSE) (Pouget et al., 2024a, b; Sohrabizadeh et al., 2021, 2023, 2022; Ferretti et al., 2018; Zhao et al., 2017; Sohrabizadeh et al., 2023) or source-to-source compilers (Lai et al., 2019; Ye et al., 2022) to guide the synthesis process. These pragmas aid in optimizing the generated hardware code. Despite the improvements in productivity, both code transformation and ensuring the correctness of the transformations remains challenging. The HLS design space encompasses various code transformations and pragma placements. However, exploring this joint space can be daunting due to its vastness, with millions of potential points, and traditional analytical methods struggle to efficiently navigate it. Key issues include the lack of convexity and regularity in this space. Moreover, separating code transformation and pragma insertion into distinct optimization problems complicates the search and may necessitate initiating the search for one optimization problem upon any changes occurring in the other.
Our primary research objective is to develop a system capable of autonomously conducting search-friendly code transformation, tile size selection to cache the data on-chip and integrating hardware pragmas for HLS to enhance parallelization. We seek to attain a favorable QoR, especially for computation-bound designs, where maximizing parallelism is crucial for optimizing QoR. To address this challenge, we introduce Sisyphus, a framework built on top of HLS compilers. This framework automates code transformation, including loop splitting, permutation, tiling, and pragma insertion for unrolling, pipelining, and on-chip data caching while partitioning arrays to ensure efficient parallelization, all within a single optimization problem. Even though several exploration methods could be used, we chose to employ a Non-Linear Programming (NLP) approach. This method facilitates rapid exploration of the entire theoretical space. We develop an analytical model that integrates considerations of latency and resource utilization, building upon previous research (Pouget et al., 2024b, a). Our focus lies specifically on regular loop-based kernels (PolyBench, [n. d.]), ensuring meticulous control over the correctness of code transformations and the accuracy of cost models. This model relies on parameters derived from both the program’s schedule and the pragma configuration. To facilitate seamless code transformation while respecting constraints and pragma insertion, we have designed a novel template tailored to these objectives. In our template, we incorporate a two-level tiling strategy, where each tiling level corresponds to particular HLS optimizations such as fine-grained unrolling, pipelining, coarse-grained unrolling, or tiling. As a result, the loop trip counts become variables, with each loop associated with a specific pragma determined by the tiling level. By solving the NLP problem, we determine these trip counts, as well as other parameters like the array partitioning factor, enabling the automatic generation of the corresponding C++ code.
In summary, we introduce the following contributions:
  • A novel template that streamlines code transformation, pragma insertion, and tiles size selection for data caching by consolidating these tasks into a single optimization problem. This approach not only simplifies the search space but also ensures that only legal transformations are considered.
  • We develop a novel NLP approach tailored to exploring this joint space, particularly focusing on regular loop-based kernels. The NLP allows for the exploration of a space whose resources are set by the user.
  • Our framework acts as a comprehensive, fully automated system, offering end-to-end functionality. With it, we conduct thorough evaluations and attain QoR designs that are comparable or superior to those achieved by AutoDSE, NLP-DSE and HARP, all without the necessity of Design Space Exploration. Furthermore, Sisyphus accurately identifies the appropriate transformation, even in cases where no transformation is required.
The paper is organized into the following sections. Section 2 lays out the motivation behind our approach and the proposed solution. Following that, Section 3 presents the range of code transformation and pragma insertion we consider. In Section 4, we introduce a non-linear formulation based on this model to automatically discover schedule and pragma configurations through NLP optimization. Moving forward, Section 5 elaborates on our code generation process. Finally, sections 6, 7 and 8 are dedicated to evaluating our method, presenting related work, and drawing conclusions.
2. Background and Motivation
2.1. Space to explore
The exploration space for optimizing an HLS design encompasses pragma insertion, code transformation, and data caching, all while adhering to resource constraints. While various objective functions can be considered, our focus in this work is on minimizing latency by maximizing parallelism. This objective function is especially potent for computation-bound kernels, addressing the bottleneck caused by insufficient parallelization. These different elements of the space (pragma insertion, code transformation, and data caching) are inherently interconnected. Inserting unrolled pragmas influences array partitioning and thus the utilization of BRAM. The schedule and pragmas influence the possibility of caching data on-chip, and the choice of schedule impacts how we want to unroll based on loop properties (e.g., reduction loop), and so on. Thus, the choice of one element affects the others, necessitating backtracking if the spaces are separated. This complexity complicates the exploration of the space and also requires constant verification of the legality of transformations.
2.2. Limitation of the current DSE
Several frameworks (Lai et al., 2019; Ye et al., 2022; Zhao et al., 2022) enable code transformations and pragma insertion, but their spaces are limited for both. Transformations are restricted to loop property-based permutations or the use of Pluto (Bondhugula et al., 2008), which tiles the program but may restrict parallelization for FPGAs and does not consider FPGA-specific optimizations such as pipelining or array partitioning.
On the other hand, numerous Design Space Explorations (DSEs) (Sohrabizadeh et al., 2021, 2023, 2022; Pouget et al., 2024b) explore pragma insertion for a fixed loop order. Although the user may initially set the schedule, it can be difficult, even for experts, to predict which transformations are necessary prior to pragma insertion. Furthermore, the multitude of code transformations renders the exploration of each pragma impractical in terms of scalability.
We will now examine examples that highlight the limitations of current methods. For this, we use two different HLS DSE methods, AutoDSE (Sohrabizadeh et al., 2021) and NLP-DSE (Pouget et al., 2024b, a). A comparison with HARP (Sohrabizadeh et al., 2023) will be presented in Section 6. AutoDSE trents the compiler Merlin (Xilinx, 2024) as a black box, adjusting pragmas based on identified bottlenecks from previous iterations. In contrast, NLP-DSE employs Nonlinear Programming DSE, utilizing a lower bound-based objective function to achieve high QoR within a short timeframe. Both of this DSE use the AMD/Xilinx source-to-source compiler Merlin (Xilinx, 2024). The compiler employs different code transformations such as strip-mining via the TILE pragma or for partial loop unrolling. It typically avoids permutations, except when partially unrolling the two innermost loops, in which case the compiler strip-mines these loops and applies permutations to the resulting fully unrolled ones.
In the upcoming examples, we utilize the HLS compiler Vitis 2023.2, excluding the unsafe-math option. We examine two scenarios: the kernel gemm from Polybench (Polybench, [n. d.]) as shown in Listing 1, and CNN.
1for (i = 0; i < 200; i++) {
2 for (j = 0; j < 220; j++)
3 C[i][j] *= beta;
4 for (k = 0; k < 240; k++)
5 for (j = 0; j < 220; j++)
6 C[i][j] += alpha * A[i][k] * B[k][j];}
Listing 1: Code original of gemm
The gemm kernel’s original loop order, as illustrated in Listing 1, situates a non-reduction loop innermost for the second statement. This arrangement facilitates AutoDSE and NLP-DSE in unrolling this loop and pipelining the reduction loop. However, partially unrolling the reduction loop in this scenario diminishes performance due to the increased initiation interval (II) of the pipeline resulting from unrolling the reduction loop. As a consequence, the designs yielding the best QoR for these two methods achieve a throughput of approximately 20 GF/s. Listing 1 highlights the design found by AutoDSE. The pragmas with $factor=1$ or pipeline off are not utilized in the final design. However, they illustrate the potential space explored by AutoDSE with various factors. In such instances, a transformation becomes imperative to augment parallelization.
1#pragma ACCEL parallel factor=1
2#pragma ACCEL pipeline off
3#pragma ACCEL TILE FACTOR=5
4for (i = 0; i < 200; i++) {
5#pragma ACCEL tile factor=1
6#pragma ACCEL pipeline off
7#pragma ACCEL parallel FACTOR=4
8 for (j = 0; j < 220; j++)
9 C[i][j] *= beta;
10#pragma ACCEL tile factor=1
11#pragma ACCEL parallel factor=1
12#pragma ACCEL pipeline flatten
13 for (k = 0; k < 240; k++)
14#pragma ACCEL pipeline off
15#pragma ACCEL tile factor=1
16 for (j = 0; j < 220; j++) // fully unrolled
17 C[i][j] += alpha * A[i][k] * B[k][j];}
Listing 2: Code of gemm found by AutoDSE
When dealing with Convolutional Neural Networks (CNNs), there’s significant room for code optimization, especially considering the need for tiling. Each loop split introduces various permutations, leading to a vast range of options, totaling $2.23\times 10^{13}$ potential combinations of loop orders and tile sizes. Despite achieving satisfactory QoR — 42.15 GF/s for AutoDSE and 31.80 GF/s for NLP-DSE — both methods fall short of fully exploiting the available level of parallelism, hinting at further optimization potential. However, exhaustively exploring each code transformation proves practically infeasible due to the sheer volume of possibilities. While the Merlin compiler can leverage strip mining to enhance on-chip data caching potential, its capabilities in this regard are constrained by the inability to permute the loops.
2.3. Overview of Sisyphus
To address these limitations, we propose Sisyphus, which focuses on linear code to offer a space capable of exploring only legal code transformations, pragma insertions, and tile size selections as a single optimization problem. Leveraging a Non-Linear Programming (NLP) approach, Sisyphus can efficiently explore this space, benefiting from accurate modeling enabled by compile-time analysis feasible for linear code. The space comprises three levels of tiling, each corresponding to a specific optimization: fine-grained unrolling, pipelining, and coarse-grained unrolling, with loops executed sequentially and data transferred between off-chip and on-chip.
For gemm, Sisyphus excels in boosting parallelism by splitting and permuting loops, all without depending on the reduction loop. In this particular setup, the NLP discovers the configuration showcased in Listing 3.
1// possibilities to cache on-chip A, B and C
2/********* Level 0 *********/
3for (int i0 = 0; i0 < 1; i0++)
4// possibilities to cache on-chip C
5 for (int j0 = 0; j0 < 1; j0++)
6// possibilities to cache on-chip C
7/********* Level 1 *********/
8 for (int i1 = 0; i1 < 1; i1++)
9 for (int j1 = 0; j1 < 55; j1++)
10#pragma HLS pipeline
11/********* Level 2 *********/
12 for (int i2 = 0; i2 < 200; i2++)
13#pragma HLS unroll
14 for (int j2 = 0; j2 < 4; j2++)
15#pragma HLS unroll
16 C[i0*200+i2][j0*220+j1*4+j2] *=beta ;
17/********* Level 0 *********/
18for (int i0 = 0; i0 < 1; i0++)
19// possibilities to cache on-chip A and C
20 for (int j0 = 0; j0 < 1; j0++)
21// possibilities to cache on-chip B and C
22 for (int k0 = 0; k0 < 60; k0++)
23// possibilities to cache on-chip A and B
24/********* Level 1 *********/
25 for (int i1 = 0; i1 < 1; i1++)
26 for (int k1 = 0; k1 < 1; k1++)
27 for (int j1 = 0; j1 < 220; j1++)
28#pragma HLS pipeline
29/********* Level 2 *********/
30 for (int i2 = 0; i2 < 200; i2++)
31#pragma HLS unroll
32 for (int j2 = 0; j2 < 1; j2++)
33#pragma HLS unroll
34 for (int k2 = 0; k2 < 4; k2++)
35#pragma HLS unroll
36 i = i0*200+i2; j = j0*220+j1*1+j2;
37 k = k0 * 4 + k2;
38 C[i][j] +=alpha * A[i][k] * B[k][j];
Listing 3: Code of gemm found by Sisyphus
Both pipelined loops feature an II of 1, enabling a throughput of 210 GF/s. Partially unrolling the reduction loop allows for parallel execution of the two multiplications.
Concerning CNN, the NLP swiftly identifies a solution within 4.5 minutes. This solution optimizes tile configuration, caches a data tile of the array on-chip, partially unrolls one reduction loop and two non-reduction loops, and pipelines a non-reduction loop. As a result, the design achieves a throughput of 339 GF/s. The design discovered by Sisyphus is showcased in Listing 4.
1for (int i0 = 0; i0 < 16; i0++) {
2 load_weight(weight,vweight,i0); // size 16*256*5*5
3 load_output(output, voutput, i0); // size 16*224*224
4 for (int j0 = 0; j0 < 256; j0++) {
5 load_input(input, vinput, j0); // size 228*228
6 for (int q0 = 0; q0 < 5; q0++)
7 for (int w0 = 0; w0 < 14; w0++)
8 for (int h1 = 0; h1 < 224; h1++)
9#pragma HLS pipeline
10 for (int i2 = 0; i2 < 16; i2++) // unrolled
11 for (int w2 = 0; w2 < 16; w2++) // unrolled
12 for (int p2 = 0; p2 < 5; p2++){ // unrolled
13 i = ...
14 output[i2][h][w] +=weight[i2][j][p][q]
15 * input[j2][h+p][w+q];}}
16 store_output(voutput, output, i0);}
Listing 4: Code of CNN found by Sisyphus
3. Unified Space
We proceed to develop a methodology that integrates code transformation, tile size selection and pragma insertion into a unified optimization framework. This involves implementing maximal distribution for code transformation, followed by the design and implementation of a template capable of executing code transformation, tiles size selection and pragma insertion simultaneously. We illustrate our code transformation process step by step using the example provided in Listing 1.
Maximum Distribution  To kickstart the process, we begin by thoroughly distributing the program, with the goal of maximizing the distance between each statement. This distribution strategy creates ample opportunities for parallelization by maximizing the separation between dependencies. Typically, this results in one statement per loop body with perfectly nested loops. To achieve this, we employ ISCC (Verdoolaege, 2011). It allows us to explore various schedule distributions and validate the legality of transformations by ensuring dependency constraints are preserved. After distribution, the example (Listing 1) yields two loop bodies. The first statement iterates through loops $i$ and $j$, while the second statement iterates through loops $i$, $k$, and $j$.
Strip-mining  Following maximal distribution, we perform strip-mining on each loop twice. This process replaces the original loop, with a trip count $TC\_I0$, with three new loops, each having trip counts $TC\_I0\_0,TC\_I0\_1$ and $TC\_I0\_2$ with $TC\_I0\_0\times TC\_I0\_1\times TC\_I0\_2=TC\_I0$. Since strip mining is inherently legal, there is no necessity to validate the legality of this transformation. Subsequently, we can arrange the loops in different permutations if it is legal. Following the implementation of three levels of strip-mining, the first loop body of the code from Listing 1 undergoes transformation into the code depicted in Listing 5.
1for (i0 = 0; i0 < TC_I0_0; i0++)
2 for (i1 = 0; i1 < TC_I0_1; i1++)
3 for (i2 = 0; i2 < TC_I0_2; i2++)
4 for (j0 = 0; j0 < TC_J0_0; j0++)
5 for (j1 = 0; j1 < TC_J0_1; j1++)
6 for (j2 = 0; j2 < TC_J0_2; j2++)
7 i = i0 * TC_I0_1 * TC_I0_2 + i1 * TC_I0_2 + i2;
8 ...
9 C[i][j] *= beta;
Listing 5: Code of the first loop body of gemm with three level strip mining
Following the explanation provided in the next paragraph, strip mining allows us to create opportunities for fine-grained unrolling, pipelining, sequential execution, coarse-grained unrolling and tiling if loop permutation is legal. This expands the range of possibilities, allowing each loop to integrate diverse hardware directives and be reorganized across different levels of the code.
Loop Permutation  Next, we consider possible permutations of the strip-mined loops. If these permutations are legal, we establish three loop levels. We ensure legality by checking dependency preservation using ISCC (Verdoolaege, 2011).
The innermost level facilitates fine-grained unrolling (complete unrolling), which increases parallelism by duplicating statements and can utilize tree reduction if a reduction loop is unrolled and the option is enabled. The middle level facilitates pipelining, enhancing throughput by overlapping loop iterations. Finally, the outermost level coordinates coarse-grained unrolling, tiling, and/or sequential execution, optimizing to increase the parallelism, and controlling the size of the on-chip buffer through tiling. If the loop body consists of only one statement and loop permutation is feasible, we disable coarse-grained unrolling. This will achieve equivalent results because our template permits unrolling the same loop innermost for fine-grained unrolling. Furthermore, coarse-grained unrolling entails duplicating each array, including read-only arrays, to ensure full overlapping of the modules, even when there are no dependencies between the concurrently executed modules.
In scenarios where permutations enable us to achieve perfectly nested loops capable of being pipelined, only one loop can usually be pipelined. Trying to pipeline all loops simultaneously, akin to pragma loop_pipelined, significantly complicates the design and extends synthesis times unreasonably. Thus, if a loop in the middle level is pipelined, it implies that the trip counts of the other loops (in the middle level) are reduced to one, ensuring manageable design complexity and synthesis times. As explained earlier, loop order selection is not necessary at the pipeline level. In the fine-grained level, arranging loops is straightforward; placing the reduction loop innermost enables efficient tree reduction. Conversely, at the coarse-grained or tiled level, prioritizing memory transfer latency is key. Therefore, we opt for a loop order that maintains the output stationary, minimizing costs associated with loading on-chip and storing off-chip during tile iterations, especially when the output is not fully transferred to the chip.
In the context of our example, these transformations yield the code presented in Listing 3, with each level corresponding to specific transformations. It is notable that our framework maintains the original loop order. In Listing 3, if the loop on line 12 ($i2$) iterates once and the loop on line 3 ($i0$) iterates 200 times, we keep the original loop order, then proceed to pipeline and partially unroll loop $j$.
Fusion  We have opted not to incorporate fusion into our model for several reasons. Firstly, our objective is to maximize parallelization for computation-bound kernels. Fusion may reduce or minimize the dependency distances between statements, limiting potential parallelization opportunities. Additionally, since our focus is on computation-bound kernels, the primary bottleneck lies in computation. Therefore, we are willing to incur a minor cost, even if it involves transferring a tiled array multiple times.
Space  Therefore, the challenge involves determining the trip counts of each loop (e.g., $TC\_I0\_0$). Additionally, the design must adhere to resource constraints, including the usage of DSP and on-chip memory. Hence, the trip counts corresponding to fully unrolled loops must be constrained to avoid over-utilization of DSP resources. The on-chip buffers must not exceed the on-chip memory capacity. Consequently, for a large problem size, the outermost level will enable control over the buffer sizes. Thus, the space we consider encompasses the original schedule on which we can insert pragmas, but it also includes all the transformations mentioned previously for which the pragmas are inserted.
4. NLP
Presented here is a comprehensive set of constraints and variables employed in a non-linear program aimed at discovering the theoretical solution space outlined in Section 3. We employ the methodology proposed by (Pouget et al., 2024b, a), adapted to our specific context. Similar to their approach, we establish an estimation of the latency by assuming optimistic DSP utilization, assuming perfect resource reuse. However, we afford the user the flexibility to adjust the DSP limit or opt for a pessimistic DSP utilization scenario, where no reuse between statements is considered. Moreover, users have the ability to adjust the size of on-chip memory, define the maximum number of array partitions, and customize the latency and resources allocated for each operation. This flexibility ensures adaptability across various platforms and compilers.
To gather all the required information for the NLP, we utilize PoCC (PoCC, [n. d.]) to extract the intermediate representation.
4.1. Variables
Tables 1 and 2 delineate the sets, variables, and constants utilized in our NLP formulation.
Set Description
$\mathcal{L},\mathcal{A},\mathcal{S}$ the set of loops, arrays and statements
$\mathcal{S}_{1}$ the set of statements alone in a loop body
$\mathcal{O}_{s}$ the list of operations of the statements $s$
$\mathcal{L}_{s}$ the set of nested loops which iterate the statement $s$
$\mathcal{L}_{s}^{red}$ the set of reduction loops which iterate the statement $s$
$\mathcal{C}_{a_{d}}$ the set of loops which iterates the array $a$ at the dimension $d$
$\mathcal{V}$ Level of the strip mining, 0 for coarse-grained/sequential, 1 for pipeline and 2 for fine-grained unrolled
$AP_{a,d}$ Array Partition for the array $a$ in dimension $d$
Constant Description
$TC_{l}$ Trip Count of the loop $l$ before strip-mining
$II_{l}$ II of the loop $l$
$IL_{par},IL_{red}$ Iteration Latency of the operations without ($par$) and with ($red$) dependencies of the statement $s$
$DSP_{s_{op}}$ Number of DSP used for the statement $s$ for the operation $op$
$DSP_{available}$ Number of DSP available for the FPGA used
$max_{part}$ Maximum array partitioning defined by the user
$ft\_arr_{a}\_loop_{l}$ Footprint of the array $a$ if transferred to on-chip after the loop $l$
$reuse_{opt}$ Boolean for optimistic reuse
Table 1. Overview of the set and constant employed in formulating the NLP
Variable Description
$tc_{l,level}$ TC of the loop $l$ for the level of strip-mining
$loop_{l}\_UF$ Coarse-grained unroll factor of the loop $l$ at level 0 of the strip-mining
$loop_{l}\_pip$ Boolean to know if the loop $l$ is pipelined at level 1 of the strip-mining
$cache_{l}\_arr_{a}$ Boolean to know if the array $a$ is transferred on-chip after the loop $l$ at level 0 of the strip-mining
Table 2. Overview of the variable employed in formulating the NLP
4.2. Constraints
Now, we explore the precise meaning and implications of each constraint.
Trip Count  Equation 1 constrain the trip count of each loop, ensuring that the product of the trip counts equals the original trip count.
(1) $\forall l\in\mathcal{L},\prod_{v\in\mathcal{V}}TC_{l,v}=TC_{l}$
Coarse-grained unrolling  We enforce coarse-grained unrolling solely for non-reduction loops (Eq. 4), where the unroll factor (UF) must divide and be less than or equal to the trip count of the current loop (Eq. 5 and 2). As elaborated in Section 3, coarse-grained unrolling is disregarded if there is only one statement within the loop (Eq. 3).
(2) $\forall l\in\mathcal{L},loop_{l}\_UF\leq TC_{l,0}$
(3) $\forall s\in\mathcal{S}_{1},\forall l\in\mathcal{L}_{s},loop_{l}\_UF=1$
(4) $\forall s\in\mathcal{S}_{1},\forall l\in\mathcal{L}_{s}^{red},loop_{l}\_UF=1$
(5) $\forall l\in\mathcal{L},loop_{l}\_UF\%TC_{l,0}==0$
Pipeline  Constraints 6, 7, 8, and 9 facilitate the selection of loop pipelining, allowing only one pipeline per nested loop and computing the initiation interval (II) in accordance with the chosen pipelined loop. The II is determined based on the loop’s properties and the iteration latency of the reduction operation, specifically when the loop being pipelined is a reduction loop, following the approach outlined in (Pouget et al., 2024b, a).
(6) $\forall l\in\mathcal{L},loop_{l}\_pip\in\{0,1\}$
(7) $\forall l\in\mathcal{L},(1-loop_{l}\_pip)\times TC_{l,1}==1$
(8) $\forall s\in\mathcal{S},\sum_{l\in\mathcal{L}_{s}}loop_{l}\_pip\leq 1$
(9) $\forall s\in\mathcal{S},II_{s}=\sum_{l\in\mathcal{L}_{s}}loop_{l}\_pip\times II_{l}$
On-chip memory  Constraints 10 and 11 ensure that each array is cached on-chip at a single location within the code, and that the on-chip tiled data fits within the available on-chip memory.
(10) $\forall l\in\mathcal{L},\forall a\in\mathcal{A},cache_{l}\_arr_{a}\in\{0,1\}$
(11) $\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}cache_{l}\_arr_{a}\times ft\_array_{a}\_loop_{l}\leq Mem$
Array partitioning  Equations 12, 13 and 14 guarantee that the maximum array partitioning is not exceeded, and it ensures that the partitioning of the array is greater than or equal to the maximum unroll factor applied to that dimension. Thanks to Equation 13, which ensures that each unroll factor ($TC_{l,2}$) divides $AP_{a,d}$, $AP_{a,d}$ directly determines the array partitioning applied in the generated code.
(12) $\forall a\in\mathcal{A},\prod_{d\in\mathbb{N}}AP_{a,d}\leq max_{part}$
(13) $\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}\%TC_{l,2}==0$
(14) $\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}\geq TC_{l,2}$
DSP utilization  Lastly, Equations 15, 16, 17, and 18 compute the DSP utilization under optimistic (full reuse) and pessimistic (no reuse) DSP reuse assumptions.
(15) $DSPs\_used_{opt}=\sum_{op\in\{+,-,*,/\}}\max_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})$
(16) $DSPs\_used_{pes}=\sum_{op\in\{+,-,*,/\}}\sum_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})$
(17) $reuse_{opt}\times DSPs\_used_{opt}\leq DSP_{available}$
(18) $(1-reuse_{opt})\times DSPs\_used_{pes}\leq DSP_{available}$
4.3. Objective Function
We structure the objective function akin to the approach described in (Pouget et al., 2024b, a), tailoring it to meet the specific needs of our problem. Here, $Lat_{2}$ corresponds to the fine-grained unrolled tile, $Lat_{1}$ denotes the pipeline tile that incorporates the fine-grained unrolled tile, and recursively, $Lat_{0}$ encompasses $Lat_{1}$, enabling coarse-grained unrolling and facilitating on-chip data caching.
$\left.\begin{cases}&Lat_{2}=IL_{par}+IL_{seq}\times\prod_{l\in\mathcal{L}^{red}}\times\log_{2}(TC_{l,2})\\ &Lat_{1}=Lat_{2}+II\times(TC_{l,1}-1)\\ &Lat_{0}=\prod_{l\in\mathcal{L}}\frac{TC_{l,0}}{loop_{l}\_UF}\times Lat_{1}\\ &L_{mem}=\sum_{l\in\mathcal{L}}\max_{a\in\mathcal{A}}(cache_{l}\_arr_{a}\times ft\_array_{a}\_loop_{l})\\ &obj\_func=Lat_{0}+L_{mem}\end{cases}\right.$
5. Code Generation and Optimization
The NLP allows us to find the trip count of each loop, the array partitioning for each array, and the size of the on-chip buffers. By considering the loop order and the pragmas applied to each loop as described in Section 3, we can directly generate the code. We automatically generate the functions that load data from off-chip to on-chip with the maximum burst size possible and vice versa with the functions that store the data off-chip. Additionally, we incorporate further optimizations to enhance the QoR.
5.1. Optimization for non-constant trip count
To optimize loops with non-constant trip counts, we implement a code transformation that preserves the NLP’s estimation while simplifying compilation for the HLS compiler. We achieve this by replacing loops with non-constant trip counts with the maximal trip count computed using PoCC (PoCC, [n. d.]). Subsequently, we create a function for all statements iterated by this loop, ensuring compliance with the constraints of the non-constant trip count loop. This function returns the computation if the constraints are met or returns just the value of the output otherwise. Introducing a condition solely above the statement results in excessive compilation times. Therefore, we employ these techniques to reduce compilation time and achieve designs with a good QoR.
5.2. Overlapping computation and communication
The constraints of the NLP consider the sum of the footprints of all arrays used during the design execution. Thus, we have the option, while already anticipating resource constraints, to overlap the transfer of arrays between off-chip and on-chip (or vice versa) at the same level or between the transfer of the arrays needed for the next loop body with the computation of the current loop body.
To achieve this, we separate each loop body into different functions, as well as the functions that transfer input data between off-chip and on-chip. Additionally, we create new functions that execute these independent functions in parallel. To ensure parallel execution, within these newly created functions designed for overlap, we include only independent functions.
5.3. Code transformation for HLS
To simplify the understanding of the HLS compiler, particularly to ensure that the compiler can pipeline loops with the correct initiation interval (II), we simplify the reductions (regardless of whether we use tree reductions) in the fine-grained unrolled part. As specified in section 3, the fully unrolled loops are placed innermost. Therefore, before the reduction, we create a variable to accumulate the reduction, which we then add to the output.
Additionally, we incorporate the pragma loop_flatten off on all reduction loops above the pipeline. While theoretically, this adjustment could enhance the QoR, in practice, the compiler might choose not to pipeline at all if it cannot flatten the loop and pipeline correctly.
6. Evaluation
6.1. Setup
We employ kernels sourced from Polybench/C 4.2.1 (PolyBench, [n. d.]), supplemented by a Convolutional Neural Network (CNN) kernel and matrix multiplication tasks akin to those in the BERT transformer model. Our computations utilize single-precision floating-point data types as the default, facilitating comparison with AutoDSE and NLP-DSE. We operate on medium-sized datasets from PolyBench/C (PolyBench, [n. d.]). We also compare our approach with HARP (Sohrabizadeh et al., 2023), which seeks to enhance High-Level Synthesis design exploration by integrating a Graph Neural Network model that predicts HLS tool behavior. They conduct exploration within this space utilizing the GNN model for one hour and then synthesize the top-10 designs. We use the medium-sized Polybench kernels found in their training set with double-precision floating-point data types, aiming to deploy HARP under optimal conditions. As detailed in (Pouget et al., 2024b, a), HARP’s effectiveness is not universally applicable without the necessary fine-tuning or training, particularly when confronted with diverse problem sizes.
We chose these problem sizes to demonstrate the efficacy of our approach in transforming codes and inserting pragmas in scenarios where fully unrolling is not feasible. Additionally, we aim to illustrate the effectiveness of our framework in tiling code for substantial tasks such as CNN and bert_3072_100. This approach ensures that only a portion (a tile) of each array fits on-chip, aligning with the available memory constraints. The problem size and original loop order of CNN are I,J=256, H,W=224 and P,Q=5. For bert_n_m matrix multiplication the schedule and problem size are I=n, J=m, K=n (reduction).
We utilize the AMD/Xilinx Vitis HLS compiler (Xilinx, 0232) to showcase the effectiveness of Sisyphus. Report generation is performed using AMD/Xilinx Vitis 2023.2. When enabling tree reduction, we utilize the ”funsafe math optimizations” option to enable commutative/associative reduction operators, thereby facilitating reductions to be implemented in logarithmic time. In our comparison with HARP (Sohrabizadeh et al., 2023), we utilized the same parameters as the authors, employing Vitis 2021.1 and activating the option for tree reduction. As our designated hardware platform, we deploy the Xilinx Alveo U200 device with a target frequency of 250 MHz. We conduct an analysis of the kernels and automatically generate each NLP problem using PoCC (PoCC, [n. d.]). Solving these NLP problems, we employ the AMPL description language and utilize the commercial BARON solver version 21.1.13 (Tawarmalani and Sahinidis, 2005; Sahinidis, 2017). For our experimental setup, we harness the computing power of 2 AMD EPYC 7V13 64-Core Processor.
6.2. Experimental Evaluation
The aim of this evaluation is to demonstrate our capability to perform the proper transformation while simultaneously inserting the correct hardware directives. To achieve this, we compare our approach with three frameworks, that conduct design space exploration to insert pragmas with a fixed schedule.
The input codes provided are the original codes from Polybench or as described in the subsection 6.1. However, we modify the kernels containing loops with non-constant trip counts using the method outlined in Section 5.3. This adjustment ensures a fair evaluation with the transformations we apply.
We compare our method with AutoDSE (Sohrabizadeh et al., 2021) and NLP-DSE (Pouget et al., 2024a, b) described in Section 2, and HARP (Sohrabizadeh et al., 2023).
We generate the AutoDSE space automatically by running the command ds_generator. We replace the unroll factor and tile size with all factors of the trip count to ensure consistency within the space. However, since maximum array partitioning is not a parameter within AutoDSE’s space, this space, as observed later with the kernel atax, is slightly larger. AutoDSE is executed using the bottleneck method with a DSE timeout set to 1,000 minutes and a timeout of 180 minutes for each HLS synthesis. NLP-DSE is executed with the parameters specified in the paper, including a 30-minute timeout for the BARON solver and the space provided as input is $\{\infty,2048,1024,512,256,128,64,32,16,8,1\}$. For HARP, we adhere to the parameters outlined by the authors mentioned earlier.
In evaluating Sisyphus, we set a solver timeout of 14,400 seconds (4 hours) for the Baron solver, which generates only one design. By default, we assume that the compiler can efficiently reuse resources between statements that are not executed in parallel. If the synthesis results in the design exceeding DSP utilization, we relaunch the NLP with a pessimistic DSP utilization assumption—meaning we consider no reuse between statements—and regenerate the design accordingly. For heat-3d, symm, syrk, syr2k, and bert_3072_100 we had to apply the constraint indicating no reuse (Eq. 16), as the constraint with optimistic reuse (Eq. 15) resulted in kernels with resource over-utilization. The maximum partitioning constraint, $max_{part}$ (cf. Section 4), is set at 1024, as determined by our use of AMD/Xilinx compilers and FPGAs, adhering to their specified limit.
Each synthesis process begins with a C-simulation to verify the correctness of the generated code.
6.3. Comparison of the methods
6.3.1. AutoDSE and NLP-DSE
Table 3 provides a comparison among AutoDSE, NLP-DSE, and Sisyphus. The TR column denotes the status of the tree reduction option, while the T column indicates the specific transformations applied, with D representing distribution, T for data tiling, and P for permutation. Across all evaluated kernels, we consistently achieve comparable or superior performance, with the exception of the atax kernel, where AutoDSE yields designs approximately 1% to 2% faster, and heat-3d, where NLP-DSE makes a 1% improvement. However, the designs obtained by AutoDSE fall outside our defined space due to the stringent constraint of a maximum array partitioning limit of 1024 imposed by AMD/Xilinx Vitis. The design found by NLP-DSE for heat-3d shares the same schedule and pragma as ours, but slight enhancements in QoR are attributed to how Merlin handles memory transfer.
{adjustwidth}
-1in-1in
GF/s Perf. Improvement
Kernel TR T Sisyphus AutoDSE NLP-DSE
atax 1 D 1.96 0.99x 1.00x
atax 0 D 1.86 0.98x 1.00x
heat-3d - - 131.33 35.09x 0.99x
jacobi-2d - - 128.83 13.06x 6.60x
2mm 1 D,P 164.34 403.55x 1.33x
2mm 0 D,P 162.24 775.26x 1.36x
bert_100_64 1 D,P 88.30 1.08x 1.03x
bert_100_64 0 D,P 76.76 1.09x 1.09x
bert_100_768 1 D,P 218.08 1.04x 1.10x
bert_100_768 0 D,P 216.95 1.13x 1.13x
bert_100_3072 1 D,P 220.29 1.10x 1.10x
bert_100_3072 0 D,P 219.97 1.11x 1.11x
gemm 1 D,P 198.26 1.79x 1.51x
gemm 0 D,P 210.63 9.25x 9.57x
bicg 1 D,P 1.96 1.97x 1.97x
bicg 0 D,P 1.86 1.92x 1.93x
gemver 1 P 17.85 4.64x 1.77x
gemver 0 P 12.45 30.91x 7.24x
mvt 1 P 13.71 1.76x 1.76x
mvt 0 P 10.00 1.45x 1.45x
bert_3072_100 1 D,P,T 42.93 2.98x 3.31x
bert_3072_100 0 D,P,T 44.67 536.42x 626.60x
cnn 1 D,P,T 340.20 8.00x 8.91x
cnn 0 D,P,T 339.34 8.05x 10.67x
doitgen 1 D,P 53.22 1.33x 2.66x
doitgen 0 D,P 54.05 5.46x 2.71x
symm 1 D,P 236.82 8.15x 7.20x
symm 0 D,P 236.56 8.14x 7.19x
syr2k 1 D,P 428.12 9.28x 3.12x
syr2k 0 D,P 428.39 17.57x 3.13x
syrk 1 D,P 306.07 12.44x 4.40x
syrk 0 D,P 308.71 12.55x 4.43x
Average 153.65 59.99x 22.85x
Geo Mean 69.09 5.97x 2.89x
Table 3. Comparison of the Throughput (GF/s) achieved with Sisyphus, AutoDSE and NLP-DSE
Sisyphus averages a speedup of 59.99x and 22.85x over AutoDSE and NLP-DSE, respectively, across the evaluated kernels. In terms of geometric mean, Sisyphus achieves a speedup of 5.97x and 2.89x.
Conserve the original schedule when needed  In the case of atax and jacobi-2d, Sisyphus preserves the original schedule and solely inserts pragmas. Thus, it effectively maintains the existing schedule while efficiently incorporating pragmas, a process comparable to that of two DSEs dedicated to pragma insertion. Regarding heat-3d, the original design opts for partial unrolling of the three loops with unroll factors of 2, 2, and 38, respectively, enabling a level of parallelism unattainable without code transformation. However, upon evaluation, we found that this design was over-utilizing resources. Therefore, the NLP solution, using a pessimistic constraint for DSP reuse, resulted in a design with the original schedule.
Although we note improvements for jacobi-2d, it is noteworthy that the same pragma insertion was utilized in at least one DSE. However, despite this consistency, AMD/Xilinx Vitis 2023.2 failed to synthesize the design generated by the source-to-source compiler Merlin, which the two DSEs employ.
Code transformation to manage reduction loop  In the cases of 2mm, bert, and gemm, we observe notable improvements resulting from code transformations, particularly in effectively managing the reduction loop. This enhancement is particularly evident with bert_3072_100, where the reduction loop exhibits a large trip count, as well as in gemm when tree-reduction is not used.
Code transformation with tiling  The selection of tile size becomes crucial for determining which array sizes, particularly for bert_3072_100 and CNN with their large arrays, should be cached on-chip and where they should be transferred. Organizing the outermost loop as output stationary (cf. Section 5.3), prompts the NLP to perform partial transfers of the array iterated by these loops in these examples. Consequently, the NLP’s choices in these scenarios aim to optimize data reuse and minimize redundant memory transfers. However, for CNN, its large size requires multiple transfers of the array input (cf. Listing 4).
Code transformation for memory-bound kernel  For mvt and gemver, both memory-bound kernels, effective optimization entails loop permutations to prevent redundant array transfers and accommodate array partitioning within the constraints of the AMD/Xilinx compiler’s 1024 limit. Sisyphus adeptly manages these loop permutations and pragma insertions, guaranteeing optimized performance for these kernels. Additionally, in the case of gemver without tree reduction, the NLP reverses all loop orders to pipeline the non-reduction loop, further enhancing the QoR. For bicg, fully distributing and arranging the loops appropriately can enhance parallelization capabilities.
Allow to achieve a parallelism not achievable without loop transformation  For doitgen, symm, syrk, and syr2k, the original schedule imposes constraints on the achievable level of parallelism. This limitation elucidates the performance enhancements facilitated by Sisyphus. In the case of syrk and syr2k, the primary advantageous transformation is the maximal loop distribution. Meanwhile, for doitgen and symm, we observe QoR improvements because code transformations enable a combination of loop unrolling not attainable even with maximal loop distribution.
6.3.2. HARP
In Table 4 we compare the throughput and resource utilization (BRAM, DSP, LUT, FF) achieved with HARP (Sohrabizadeh et al., 2023). The enhanced performance seen in both bicg and gemver can be attributed to the same underlying factor observed in the comparison with AutoDSE and NLP-DSE. Regarding mvt, HARP produced comparable performance to our own due to their use of the Merlin source-to-source compiler, which facilitates loop permutations when partially unrolling consecutive loops. This allows for them to find the same schedule that our framework identified. However, for Gemm, the design discovered by HARP falls outside our specified parameter space, particularly because of our restriction on maximal array partitioning. Nonetheless, it is worth noting that it is the only design among the top 10 evaluated that demonstrates a performance enhancement over our framework.
GF/s Resource Utilization (%) Perf. Imp
BRAM,DSP,LUT,FF
Kernel HARP Sisyphus HARP Sisyphus
atax 1.72 1.77 78,52,49,50 38,26,25,31 1.03x
bicg 0.92 1.77 75,25,35,36 38,26,24,26 1.92x
gemm 125.59 99.96 29,80,55,40 1,67,51,34 0.80x
gemver 1.66 7.38 29,28,35,19 43,27,39,47 4.45x
mvt 7.07 7.06 40,78,43,30 21,64,33,26 1.00x
Average 27.39 23.58 1.84x
Geo Mean 4.71 6.95 1.47x
Table 4. Comparison of GF/s and resources utilization with HARP
6.4. Latency estimation
Implementing the post-optimization strategy outlined in Section 5.2 leads to a loss of the lower bound discussed in (Pouget et al., 2024a, b), primarily because we do not incorporate computation-communication overlap modeling into the NLP. This omission is deliberate, as including such modeling would significantly increase search time.
Similarly, we opt against integrating the loop_flatten pragma into our model, despite its potential automatic invocation by the compiler for loops without reductions. As detailed in Section 5, we intentionally deactivate it for loops with reductions. Although this optimization’s automatic application results in the loss of the lower bound, we have made this choice owing to its rare implementation. Our aim is to avoid reliance on an optimization that may not be consistently available.
In Figure 1, we depict the estimated latency by the NLP of the evaluated designs and juxtapose them with the latencies reported by HLS. Our post-optimization efforts reveal that most latencies estimated by NLP exceed those reported by HLS, with the exception of CNN with and without tree reduction. In the case of CNN (the last two dots), the pragmas are implemented as anticipated, but the lower bound of the computed unrolled part is not sufficiently tight.
Figure 1. Comparison of the latency reported by HLS with the estimated latency obtained through NLP
6.5. Scalability of the NLP
Among the 42 NLP evaluations (including those with pessimistic resource reuse), only 3 timed out after 4 hours. These timeouts were specifically encountered during the bicg and atax kernels with tree reduction, and the gemver kernel without tree reduction. Throughout the search process, the NLP solver BARON maintains two bounds: a lower bound, which may be unattainable, and the lowest achievable bound discovered thus far. If the solver completes before reaching the timeout, these two bounds are identical. In the cases of atax and bicg, these two bounds were within 1% of each other, while for gemver, the difference was 20%. The average solving time was 33 minutes, while the geometric mean stood at 2.94 minutes. Out of 42 instances, 35 finished within an hour.
7. Related Work
Design Space Exploration (DSE) methodologies for pragma insertion, such as those mentioned in (Pouget et al., 2024a, b; Sohrabizadeh et al., 2021, 2023, 2022; Ferretti et al., 2018; Zhao et al., 2017; Sohrabizadeh et al., 2023; Zhong et al., 2016; Wu et al., 2023; Liu and Carloni, 2013; Mahapatra and Schafer, 2014; Yu et al., 2018), yield designs with a satisfactory QoR. However, they often require extensive computational time, with AutoDSE taking up to a day to complete. While users have the option to perform code transformations before conducting DSE, the process typically involves treating each code transformation and pragma insertion as separate optimization problems. This decomposition of the problem can result in a loss of QoR because predicting which transformation is necessary to achieve an optimal design is challenging. The NLP-DSE method, as explained in (Pouget et al., 2024b, a), employs a non-linear solver to integrate pragmas into the code. We build upon the methodology developed by the authors and tailor it to our specific optimization space to determine the pragmas and schedule for the kernel.
Various code transformations have been devised for CPUs (Bondhugula et al., 2008), GPUs (Verdoolaege et al., 2013), and FPGAs (Pouchet et al., 2013; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018). While code transformations for CPUs and GPUs yield remarkable results tailored to their respective architectures, they may not be inherently suitable for FPGA targets aimed at achieving high parallelism. Pluto (Bondhugula et al., 2008), considered state-of-the-art, performs code transformations including tiling to minimize dependency distance between memory accesses, thereby facilitating data reuse. However, in our scenario, reducing dependency distance could potentially constrain parallelization efforts. Regarding (Zhao et al., 2022; Zhao and Cheng, 2021), they employ Pluto on different scopes of the kernel for code transformation, yet its pragma insertion capabilities are limited, making it incomparable to our work. Conversely, (Pouchet et al., 2013; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) pursue a distinct objective from ours. (Pouchet et al., 2013) aims to minimize communication between off-chip and on-chip, which results in better QoR than ours for memory-bound kernels. The works (Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) concentrate on code transformations aimed at enhancing pipelining techniques. However, these objectives may not be suitable for computation-bound kernels requiring high levels of parallelization.
The (Ye et al., 2022; Lai et al., 2019; Huang et al., 2021; Chen et al., 2024; Ye et al., 2024) compilers undertake code transformation and pragma insertion. However, these transformations are somewhat limited, mostly encompassing heuristic modifications based on loop properties. Moreover, the scope of pragma insertion is more limited compared to our proposed approach.
The choice of tile sizes significantly impacts the QoR. In line with our methodology, (Liu et al., 2017a; Li et al., 2021) employ a cost model to determine the tile size. However, while (Liu et al., 2017a) focuses on minimizing communication overhead, our approach differs. On the other hand, (Li et al., 2021) investigates tiles size selection for Convolutional Neural Networks (CNNs) with three-level CPU caching.
8. Conclusion
This paper addresses challenges in optimizing HLS via code transformation and pragma insertion. While HLS and source-to-source compilers accelerate hardware design, achieving high-quality results often demands manual code intervention, prone to errors and time-consuming. To address this, Sisyphus is introduced, automating code transformation and pragma insertion for HLS into a single optimization problem. Using Non-Linear Programming (NLP), Sisyphus determines code transformations and pragma placements simultaneously, ensuring optimized performance within resource constraints.

### v4

@@
 Abstract.
-High-level synthesis, source-to-source compilers, and various Design Space Exploration techniques for pragma insertion have significantly improved the Quality of Results of generated designs.
-These tools offer benefits such as reduced development time and enhanced performance.
-However, achieving high-quality results often requires additional manual code transformations and tiling selections, which are typically performed separately or as pre-processing steps.
-Although DSE techniques enable code transformation upfront, the vastness of the search space often limits the exploration of all possible code transformations, making it challenging to determine which transformations are necessary.
-Additionally, ensuring correctness remains challenging, especially for complex transformations and optimizations.
-To tackle this obstacle, we first propose a comprehensive framework leveraging HLS compilers.
-Our system streamlines code transformation, pragma insertion, and tiles size selection for on-chip data caching through a unified optimization problem, aiming to enhance parallelization, particularly beneficial for computation-bound kernels.
-Them employing a novel Non-Linear Programming (NLP) approach, we simultaneously ascertain transformations, pragmas, and tile sizes, focusing on regular loop-based kernels.
-Our evaluation demonstrates that our framework adeptly identifies the appropriate transformations, including scenarios where no transformation is necessary, and inserts pragmas to achieve a favorable Quality of Results.
+High-Level Synthesis compilers and Design Space Exploration tools have greatly advanced the automation of hardware design, improving development time and performance.
+However, achieving a good Quality of Results still requires extensive manual code transformations, pragma insertion, and tile size selection, which are typically handled separately.
+The design space is too large to be fully explored by this fragmented approach.
+It is too difficult to navigate this way, limits the exploration of potential optimizations, and complicates the design generation process.
+To tackle this obstacle we propose Sisyphus, a unified framework that automates code transformation, pragma insertion, and tile size selection within a single optimization problem.
+By leveraging Nonlinear Programming, our approach efficiently explores the vast design space of regular loop-based kernels, automatically selecting loop transformations and pragmas that minimize latency.
+Evaluation against state-of-the-art frameworks, including AutoDSE, NLP-DSE, and ScaleHLS, shows that Sisyphus achieves superior Quality of Results, outperforming alternatives across multiple benchmarks.
+By integrating code transformation and pragma insertion into a unified model, Sisyphus significantly reduces design generation complexity and improves performance for FPGA-based systems.
 HLS, code transformation, pragma insertion, non-linear problem
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; October 27–31, 2024; New Jersey, USAisbn: 978-1-4503-XXXX-X/18/06
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; March xx–xx, 2025; California, USAisbn: 978-1-4503-XXXX-X/18/06
 1. Introduction
 High-Level Synthesis (HLS) compilers and source-to-source compilers are indispensable tools in accelerating hardware design by automating the translation of high-level programming languages like C/C++ or Python into hardware descriptions.
-They offer notable advantages, such as reduced development time and enhanced performance for hardware designs (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020).
-However, achieving a good quality of results (QoR) often requires manual code transformations and pragma insertions, which can be facilitated with either Design Space Exploration (DSE) (Pouget et al., 2024a, b; Sohrabizadeh et al., 2021, 2023, 2022; Ferretti et al., 2018; Zhao et al., 2017; Sohrabizadeh et al., 2023) or source-to-source compilers (Lai et al., 2019; Ye et al., 2022) to guide the synthesis process.
+They offer notable advantages, such as reduced development time and enhanced performance for hardware designs (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020; Zhang et al., 2024).
+However, achieving a good quality of results (QoR) often requires manual code transformations and pragma insertions, which can be facilitated with either Design Space Exploration (DSE) (Pouget et al., 2024a, b; Sohrabizadeh et al., 2021, 2023, 2022; Ferretti et al., 2018; Zhao et al., 2017; Sohrabizadeh et al., 2023) or source-to-source compilers (Lai et al., 2019; Ye et al., 2022, 2024; Zhang et al., 2024) to guide the synthesis process.
 These pragmas aid in optimizing the generated hardware code.
-Despite the improvements in productivity, both code transformation and ensuring the correctness of the transformations remains challenging.
-The HLS design space encompasses various code transformations and pragma placements.
-However, exploring this joint space can be daunting due to its vastness, with millions of potential points, and traditional analytical methods struggle to efficiently navigate it.
-Key issues include the lack of convexity and regularity in this space.
-Moreover, separating code transformation and pragma insertion into distinct optimization problems complicates the search and may necessitate initiating the search for one optimization problem upon any changes occurring in the other.
-Our primary research objective is to develop a system capable of autonomously conducting search-friendly code transformation, tile size selection to cache the data on-chip and integrating hardware pragmas for HLS to enhance parallelization.
+Despite their advantages, existing HLS workflows face several significant challenges.
+Previous frameworks (Ye et al., 2022, 2024; Lai et al., 2019; Zhang et al., 2024) have integrated pragma insertion and code transformation, but these efforts are somewhat limited, focusing primarily on loop permutations based on loop properties.
+Additionally, code transformations and pragma insertion are often treated as separate processes.
+Transformations such as loop tiling, splitting, and permutation are typically performed in advance or as part of a pre-processing stage, while pragmas for optimizations like loop unrolling, pipelining, and array partitioning are applied later.
+This separation introduces inefficiencies, as transformations can affect the effectiveness of pragmas, and vice versa.
+Moreover, the immense design space—encompassing millions of potential transformation and pragma combinations—makes exhaustive exploration unfeasible.
+Ensuring the correctness of these transformations while adhering to resource constraints further complicates the process.
+Our primary research objective is to develop a system capable of autonomously conducting code transformation, tile size selection to cache the data on-chip and integrating hardware pragmas for HLS to enhance parallelization.
 We seek to attain a favorable QoR, especially for computation-bound designs, where maximizing parallelism is crucial for optimizing QoR.
 To address this challenge, we introduce Sisyphus, a framework built on top of HLS compilers.
-This framework automates code transformation, including loop splitting, permutation, tiling, and pragma insertion for unrolling, pipelining, and on-chip data caching while partitioning arrays to ensure efficient parallelization, all within a single optimization problem.
-Even though several exploration methods could be used, we chose to employ a Non-Linear Programming (NLP) approach.
+This framework automates code transformation, including loop splitting, permutation, tiling, and pragma insertion for unrolling, pipelining, and on-chip data caching while partitioning arrays to ensure efficient parallelization, all within a single optimization framework.
+Even though several exploration methods could be used, we chose to employ a Nonlinear Programming (NLP) approach.
 This method facilitates rapid exploration of the entire theoretical space.
-We develop an analytical model that integrates considerations of latency and resource utilization, building upon previous research (Pouget et al., 2024b, a).
+We have developed an analytical model that integrates considerations of latency and resource utilization, building upon previous research (Pouget et al., 2024b, a).
 Our focus lies specifically on regular loop-based kernels (PolyBench, [n.
@@
 In summary, we introduce the following contributions:
-  • A novel template that streamlines code transformation, pragma insertion, and tiles size selection for data caching by consolidating these tasks into a single optimization problem.
+  • A novel template that integrates code transformation, pragma insertion, and tiles size selection for data caching by consolidating these tasks into a single optimization problem.
 This approach not only simplifies the search space but also ensures that only legal transformations are considered.
-  • We develop a novel NLP approach tailored to exploring this joint space, particularly focusing on regular loop-based kernels.
-The NLP allows for the exploration of a space whose resources are set by the user.
+  • A novel NLP-based approach specifically designed to explore the joint design space of regular loop-based kernels.
+Unlike traditional methods that involve time-consuming DSE, our approach provides solutions within seconds or minutes.
+Additionally, the NLP is user-friendly, allowing users to easily configure parameters, such as on-chip buffer size, to meet the specific requirements of the kernel.
   • Our framework acts as a comprehensive, fully automated system, offering end-to-end functionality.
-With it, we conduct thorough evaluations and attain QoR designs that are comparable or superior to those achieved by AutoDSE, NLP-DSE and HARP, all without the necessity of Design Space Exploration.
+With it, we can conduct thorough evaluations and attain QoR designs that are comparable or superior to those achieved by AutoDSE, NLP-DSE, ScaleHLS and HARP, all without the necessity of DSE.
 Furthermore, Sisyphus accurately identifies the appropriate transformation, even in cases where no transformation is required.
@@
 Following that, Section 3 presents the range of code transformation and pragma insertion we consider.
-In Section 4, we introduce a non-linear formulation based on this model to automatically discover schedule and pragma configurations through NLP optimization.
-Moving forward, Section 5 elaborates on our code generation process.
-Finally, sections 6, 7 and 8 are dedicated to evaluating our method, presenting related work, and drawing conclusions.
+In Section 4, we introduce a nonlinear formulation based on this model to automatically discover schedule and pragma configurations through NLP optimization.
+Section 5 elaborates on our post-optimizations.
+Finally, sections 6, 7, 8 and 9 are dedicated to evaluating our method, presenting related work, addressing limitations and drawing conclusions.
 2. Background and Motivation
 2.1. Space to Explore
-The exploration space for optimizing an HLS design encompasses pragma insertion, code transformation, and data caching, all while adhering to resource constraints.
-While various objective functions can be considered, our focus in this work is on minimizing latency by maximizing parallelism.
-This objective function is especially potent for computation-bound kernels, addressing the bottleneck caused by insufficient parallelization.
-These different elements of the space (pragma insertion, code transformation, and data caching) are inherently interconnected.
-Inserting unrolled pragmas influences array partitioning and thus the utilization of BRAM.
-The schedule and pragmas influence the possibility of caching data on-chip, and the choice of schedule impacts how we want to unroll based on loop properties (e.g., reduction loop), and so on.
-Thus, the choice of one element affects the others, necessitating backtracking if the spaces are separated.
-This complexity complicates the exploration of the space and also requires constant verification of the legality of transformations.
+The exploration space for optimizing an HLS design includes pragma insertion, code transformations, and data caching, all within the limits of resource constraints.
+Although different objective functions can be explored, this work specifically focuses on minimizing latency while respecting resource constraints.
+This objective function is especially potent for computation bound kernels because it addresses the bottleneck caused by insufficient parallelization.
+The various components of the exploration space—pragma insertion, code transformations, and data caching—are inherently interdependent.
+For example, unrolling pragmas affect array partitioning, which in turn influences BRAM utilization.
+Similarly, the scheduling and pragmas impact the potential for on-chip data caching, while the choice of schedule dictates how loops should be unrolled based on their characteristics, such as whether they are reduction loops.
+A reduction loop is a loop where an operation, like addition or multiplication, accumulates values across iterations to produce a single output, often by reducing a multi-element array to a scalar (e.g., loop L2 in Listing 1).
+The choice of schedule for reduction loops, as opposed to non-reduction loops (e.g., L0, L1, and L3 in Listing 1), influences unrolling strategies and overall optimization because of the dependencies inherent to the loop’s characteristics.
+Consequently, selecting one element impacts the others, often requiring backtracking when the spaces are treated separately.
+This interconnection adds complexity to the exploration process and necessitates ongoing verification to ensure the legality of transformations.
+1for (i = 0; i < 200; i++) { // L0
+2 for (j = 0; j < 220; j++) // L1
+3 C[i][j] *= beta; // S0
+4 for (k = 0; k < 240; k++) // L2
+5 for (j = 0; j < 220; j++)// L3
+[…]
+[…]
+[…]
+9for (i = 0; i < 200; i++) // L0
+10 for (j = 0; j < 220; j++) // L1
+11 C[i][j] *= beta; // S0
+12for (i = 0; i < 200; i++) // L2
+13 for (k = 0; k < 240; k++) // L3
+14 for (j = 0; j < 220; j++) // L4
+[…]
+[…]
+[…]
+18for (i0 = 0; i0 < I0_S0; i0++) // L0
+19 for (i1 = 0; i1 < I1_S0; i1++) // L1
+20 for (i2 = 0; i2 < I2_S0; i2++) // L2
+21 for (j0 = 0; j0 < J0_S0; j0++) // L3
+22 for (j1 = 0; j1 < J1_S0; j1++) // L4
+23 for (j2 = 0; j2 < J2_S0; j2++) // L5
+24 C[i][j] *= beta; // S0
+25for (i = 0; i < I0_S1; i++) // L6
+26 for (i = 0; i < I1_S1; i++) // L7
+27 for (i = 0; i < I2_S1; i++) // L8
+28 for (k = 0; k < K0_S1; k++) // L9
+29 for (k = 0; k < K1_S1; k++) // L10
+30 for (k = 0; k < K2_S1; k++) // L11
+31 for (j = 0; j < J0_S1; j++) // L12
+32 for (j = 0; j < J1_S1; j++) // L13
+33 for (j = 0; j < J2_S1; j++) // L14
+[…]
+[…]
+[…]
+[…]
+38
+[…]
+[…]
+[…]
+42On-chip mem = 320 kB}]
+43load_C(C, C_off_chip);
+44/**** Level 0 of S0 ****/
+45for (j0 = 0; j0 < 1; j0++) // L0
+46 for (i0 = 0; i0 < 4; i0++) // L1
+47/**** Level 1 of S0 ****/
+48 for (j1 = 0; j1 < 22; j1++) // L2
+49#pragma HLS pipeline
+50/**** Level 2 of S0 ****/
+51 for (j2=0; j2<10; j2++) { // L3
+52#pragma HLS unroll
+53 for (i2=0; i2<50; i2++) // L4
+54#pragma HLS unroll
+55 i = ...;
+56 C[i][j] *=beta; } // S0
+57/**** Level 0 of S1 ****/
+58for (k0 = 0; k0 < 48; k0++) // L5
+59 load_B_S1(B, B_off_chip, k0);
+60 load_A_S1(A, A_off_chip, k0);
+61 for (j0 = 0; j0 < 1; j0++) // L6
+62 for (i0 = 0; i0 < 4; i0++) // L7
+63/**** Level 1 of S1 ****/
+64 for (j1=0; j1 < 220; j1++) // L8
+65#pragma HLS pipeline
+66/**** Level 2 of S1 ****/
+67 for (j2=0; j2<1; j2++) // L9
+68#pragma HLS unroll
+69 for (i2=0; i2<50; i2++) // L10
+70#pragma HLS unroll
+71 for(k2=0;k2<5;k2++){ // L11
+72#pragma HLS unroll
+73 i = ...;
+74 C[i][j]+=alpha*A[i][k2]*B[k2][j];}// S1
+75store_C(C, C_off_chip);
+Listing 1: Original Code
+1load_C(C, C_off_chip);
+2load_A(A, a_off_chip);
+3load_B(B, B_off_chip);
+4/**** Level 0 of S0 ****/
+5for (i0 = 0; i0 < 1; i0++) // L0
+6 for (j0 = 0; j0 < 1; j0++) // L1
+7/**** Level 1 of S0 ****/
+8 for (j1 = 0; j1 < 55; j1++) // L2
+9#pragma HLS pipeline
+10/**** Level 2 of S0 ****/
+11 for (i2 = 0; i2 < 200; i2++) // L3
+12#pragma HLS unroll
+13 for (j2=0; j2<4; j2++) { // L4
+14#pragma HLS unroll
+15 i = ...;
+16 C[i][j] *= beta;} // S0
+17/**** Level 0 of S1 ****/
+18for (i0 = 0; i0 < 1; i0++) // L5
+19 for (j0 = 0; j0 < 1; j0++) // L6
+20 for (k0 = 0; k0 < 60; k0++) // L7
+21/**** Level 1 of S1 ****/
+22 for (j1 = 0; j1 < 220; j1++) // L8
+23#pragma HLS pipeline
+24/**** Level 2 of S1 ****/
+25 for (i2=0;i2<200;i2++) // L9
+26#pragma HLS unroll
+27 for (j2=0; j2<1; j2++) // L10
+28#pragma HLS unroll
+29 for(k2=0;k2<4;k2++){ // L11
+30#pragma HLS unroll
+31 i = ...;
+32 C[i][j]+=alpha*A[i][k]*B[k][j];}//S1
+33store_C(C, C_off_chip);
+Listing 2: Result found using the NLP with the constraints:
+DSP = 6,840 On-chip mem = 7.2 MB
 2.2. Limitation of the Current DSE
-Several frameworks (Lai et al., 2019; Ye et al., 2022; Zhao et al., 2022) enable code transformations and pragma insertion, but their spaces are limited for both.
-Transformations are restricted to loop property-based permutations or the use of Pluto (Bondhugula et al., 2008), which tiles the program but may restrict parallelization for FPGAs and does not consider FPGA-specific optimizations such as pipelining or array partitioning.
-On the other hand, numerous Design Space Explorations (DSEs) (Sohrabizadeh et al., 2021, 2023, 2022; Pouget et al., 2024b) explore pragma insertion for a fixed loop order.
-Although the user may initially set the schedule, it can be difficult, even for experts, to predict which transformations are necessary prior to pragma insertion.
-Furthermore, the multitude of code transformations renders the exploration of each pragma impractical in terms of scalability.
+Several frameworks (Lai et al., 2019; Ye et al., 2022; Zhao et al., 2022; Ye et al., 2024; Pouchet et al., 2013a; Zhang et al., 2024; Chen et al., 2024) enable code transformations and pragma insertion, but have limited exploration spaces (Ye et al., 2022; Zhao et al., 2022; Ye et al., 2024; Pouchet et al., 2013a; Zhang et al., 2024) or lack automatic DSE (Lai et al., 2019; Chen et al., 2024).
+Transformations are often restricted to loop property-based permutations (Lai et al., 2019; Ye et al., 2022, 2024; Pouchet et al., 2013a; Zhang et al., 2024) or rely on Pluto (Bondhugula et al., 2008), which is optimized for CPUs but overlooks FPGA-specific techniques like pipelining and array partitioning.
+Additionally, Pluto minimizes memory transfer from off-chip to on-chip (Comm), which can limit parallelism on FPGAs, unlike our focus on reducing global latency (Lat).
+Meanwhile, other DSE frameworks (Sohrabizadeh et al., 2021, 2023, 2022; Pouget et al., 2024b) explore pragmas insertion for fixed loop orders, but predicting friendly transformations beforehand is challenging.
+Exploring every possible transformation is also impractical for large design spaces.
+The key differences between the frameworks are summarized in Table 1.
+Poly-Opt-HLS POL-SCA/ Pluto Auto-DSE/ HARP Hetero-CL / Allo NLP-DSE Scale-HLS/ POM Sisy-phus
+Tiling Limit.
+Permutation Limit. Limit. Limit.
+Pragma Insert.
+Code Trans + Pragma Insert. (unified)
+Off-chip comm. gen.
+Enum. (AI, heuristics,…) Manual
+Objective Comm Comm Lat - Lat Lat Lat
+Table 1. Comparison of different frameworks
 We will now examine examples that highlight the limitations of current methods.
 For this, we use two different HLS DSE methods, AutoDSE (Sohrabizadeh et al., 2021) and NLP-DSE (Pouget et al., 2024b, a).
-A comparison with HARP (Sohrabizadeh et al., 2023) will be presented in Section 6.
-AutoDSE trents the compiler Merlin (Xilinx, 2024) as a black box, adjusting pragmas based on identified bottlenecks from previous iterations.
+A comparison with HARP (Sohrabizadeh et al., 2023) and ScaleHLS (Ye et al., 2022) will be presented in Section 6.
+AutoDSE treats the compiler Merlin (Xilinx, 2024) as a black box, adjusting pragmas based on identified bottlenecks from previous iterations.
 In contrast, NLP-DSE employs Nonlinear Programming DSE, utilizing a lower bound-based objective function to achieve high QoR within a short timeframe.
@@
 It typically avoids permutations, except when partially unrolling the two innermost loops, in which case the compiler strip-mines these loops and applies permutations to the resulting fully unrolled ones.
-In the upcoming examples, we utilize the HLS compiler Vitis 2023.2, excluding the unsafe-math option.
-We examine two scenarios: the kernel gemm from Polybench (Polybench, [n.
-d.]) as shown in Listing 1, and CNN.
-1for (i = 0; i < 200; i++) {
-2 for (j = 0; j < 220; j++)
-3 C[i][j] *= beta;
-4 for (k = 0; k < 240; k++)
-5 for (j = 0; j < 220; j++)
-6 C[i][j] += alpha * A[i][k] * B[k][j];}
-Listing 1: Code original of gemm
-The gemm kernel’s original loop order, as illustrated in Listing 1, situates a non-reduction loop innermost for the second statement.
-This arrangement facilitates AutoDSE and NLP-DSE in unrolling this loop and pipelining the reduction loop.
-However, partially unrolling the reduction loop in this scenario diminishes performance due to the increased initiation interval (II) of the pipeline resulting from unrolling the reduction loop.
-As a consequence, the designs yielding the best QoR for these two methods achieve a throughput of approximately 20 GF/s.
-Listing 1 highlights the design found by AutoDSE.
-The pragmas with $factor=1$ or pipeline off are not utilized in the final design.
-However, they illustrate the potential space explored by AutoDSE with various factors.
+A more detailed description of the Merlin compiler can be found in (Pouget et al., 2024b).
+In the following examples, we use the HLS compiler Vitis 2023.2 with the unsafe-math option disabled.
+The targeted FPGA is the AMD/Xilinx Alveo U200.
+We evaluate two scenarios: the gemm kernel from Polybench (Polybench, [n.
+d.]), as shown in Listing 1, and a convolutional neural network (CNN) layer where the problem size and original loop order of CNN are I,J=256 (Output and Input Channels), H,W=224 (Height, Weight) and P,Q=5 (Filter Height and Weight).
+The gemm kernel’s original loop order, as illustrated in Listing 1, has a non-reduction loop at the center of the second statement (L3).
+This arrangement facilitates AutoDSE and NLP-DSE in unrolling this loop and pipelining the reduction loop (L2).
+In this scenario, if the reduction loop is pipelined and partially unrolled by a factor of $uf$, performance will degrade.
+While partial unrolling allows for parallel execution of the multiplications, it increases the pipeline depth due to the loop’s dependencies, as the $uf$ additions must still be performed sequentially.
+Consequently, when the reduction loop is pipelined, the initiation interval (II) will significantly increase, becoming a multiple of both $uf$ and the latency of the reduction operation.
+This occurs because the next iteration of the loop can only begin after the $uf$ additions are completed and the output is fully written.
+Their inability to further increase parallelism due to loop order limits the best designs to a throughput of 20 GFLOPs per second (GF/s).
 In such instances, a transformation becomes imperative to augment parallelization.
-1#pragma ACCEL parallel factor=1
-2#pragma ACCEL pipeline off
-3#pragma ACCEL TILE FACTOR=5
-4for (i = 0; i < 200; i++) {
-5#pragma ACCEL tile factor=1
-6#pragma ACCEL pipeline off
-7#pragma ACCEL parallel FACTOR=4
-8 for (j = 0; j < 220; j++)
-9 C[i][j] *= beta;
-10#pragma ACCEL tile factor=1
-11#pragma ACCEL parallel factor=1
-12#pragma ACCEL pipeline flatten
-13 for (k = 0; k < 240; k++)
-14#pragma ACCEL pipeline off
-15#pragma ACCEL tile factor=1
-16 for (j = 0; j < 220; j++) // fully unrolled
-17 C[i][j] += alpha * A[i][k] * B[k][j];}
-Listing 2: Code of gemm found by AutoDSE
-When dealing with Convolutional Neural Networks (CNNs), there’s significant room for code optimization, especially considering the need for tiling.
-Each loop split introduces various permutations, leading to a vast range of options, totaling $2.23\times 10^{13}$ potential combinations of loop orders and tile sizes.
-Despite achieving satisfactory QoR — 42.15 GF/s for AutoDSE and 31.80 GF/s for NLP-DSE — both methods fall short of fully exploiting the available level of parallelism, hinting at further optimization potential.
-However, exhaustively exploring each code transformation proves practically infeasible due to the sheer volume of possibilities.
-While the Merlin compiler can leverage strip mining to enhance on-chip data caching potential, its capabilities in this regard are constrained by the inability to permute the loops.
+A single CNN layer with 6 loops results in an enormous design space that cannot be explored exhaustively.
+Loop permutations alone represent 6!
+= 720 possibilities, and tiling is required to manage array sizes exceeding on-chip memory capacity.
+While frameworks like AutoDSE and NLP-DSE achieve satisfactory throughput—42.15 GF/s and 31.80 GF/s, respectively—they are unable to explore the full design space, which includes both code transformations and pragma insertion, within a reasonable timeframe.
+For example, running AutoDSE for all possible loop permutations would take around 500 days, and for NLP-DSE, it would take 106 days.
+Other frameworks, which apply code transformations based solely on loop properties, such as prioritizing reduction loops in the outermost positions, explore a much more restricted design space, limiting their potential for further optimization.
+This underscores the challenge of navigating such a vast design space and highlights gaps in current approaches.
 2.3. Overview of Sisyphus
-To address these limitations, we propose Sisyphus, which focuses on linear code to offer a space capable of exploring only legal code transformations, pragma insertions, and tile size selections as a single optimization problem.
-Leveraging a Non-Linear Programming (NLP) approach, Sisyphus can efficiently explore this space, benefiting from accurate modeling enabled by compile-time analysis feasible for linear code.
-The space comprises three levels of tiling, each corresponding to a specific optimization: fine-grained unrolling, pipelining, and coarse-grained unrolling, with loops executed sequentially and data transferred between off-chip and on-chip.
-For gemm, Sisyphus excels in boosting parallelism by splitting and permuting loops, all without depending on the reduction loop.
-In this particular setup, the NLP discovers the configuration showcased in Listing 3.
-1// possibilities to cache on-chip A, B and C
-2/********* Level 0 *********/
-3for (int i0 = 0; i0 < 1; i0++)
-4// possibilities to cache on-chip C
-5 for (int j0 = 0; j0 < 1; j0++)
-6// possibilities to cache on-chip C
-7/********* Level 1 *********/
-8 for (int i1 = 0; i1 < 1; i1++)
-9 for (int j1 = 0; j1 < 55; j1++)
-10#pragma HLS pipeline
-11/********* Level 2 *********/
-12 for (int i2 = 0; i2 < 200; i2++)
-13#pragma HLS unroll
-14 for (int j2 = 0; j2 < 4; j2++)
-15#pragma HLS unroll
-16 C[i0*200+i2][j0*220+j1*4+j2] *=beta ;
-17/********* Level 0 *********/
-18for (int i0 = 0; i0 < 1; i0++)
-19// possibilities to cache on-chip A and C
-20 for (int j0 = 0; j0 < 1; j0++)
-21// possibilities to cache on-chip B and C
-22 for (int k0 = 0; k0 < 60; k0++)
-23// possibilities to cache on-chip A and B
-24/********* Level 1 *********/
-25 for (int i1 = 0; i1 < 1; i1++)
-26 for (int k1 = 0; k1 < 1; k1++)
-27 for (int j1 = 0; j1 < 220; j1++)
-28#pragma HLS pipeline
-29/********* Level 2 *********/
-30 for (int i2 = 0; i2 < 200; i2++)
-31#pragma HLS unroll
-32 for (int j2 = 0; j2 < 1; j2++)
-33#pragma HLS unroll
-34 for (int k2 = 0; k2 < 4; k2++)
-35#pragma HLS unroll
-36 i = i0*200+i2; j = j0*220+j1*1+j2;
-37 k = k0 * 4 + k2;
-38 C[i][j] +=alpha * A[i][k] * B[k][j];
-Listing 3: Code of gemm found by Sisyphus
-Both pipelined loops feature an II of 1, enabling a throughput of 210 GF/s. Partially unrolling the reduction loop allows for parallel execution of the two multiplications.
-Concerning CNN, the NLP swiftly identifies a solution within 4.5 minutes.
-This solution optimizes tile configuration, caches a data tile of the array on-chip, partially unrolls one reduction loop and two non-reduction loops, and pipelines a non-reduction loop.
-As a result, the design achieves a throughput of 339 GF/s.
-The design discovered by Sisyphus is showcased in Listing 4.
-1for (int i0 = 0; i0 < 16; i0++) {
-2 load_weight(weight,vweight,i0); // size 16*256*5*5
-3 load_output(output, voutput, i0); // size 16*224*224
-4 for (int j0 = 0; j0 < 256; j0++) {
-5 load_input(input, vinput, j0); // size 228*228
-6 for (int q0 = 0; q0 < 5; q0++)
-7 for (int w0 = 0; w0 < 14; w0++)
-8 for (int h1 = 0; h1 < 224; h1++)
-9#pragma HLS pipeline
-10 for (int i2 = 0; i2 < 16; i2++) // unrolled
-11 for (int w2 = 0; w2 < 16; w2++) // unrolled
-12 for (int p2 = 0; p2 < 5; p2++){ // unrolled
-13 i = ...
-14 output[i2][h][w] +=weight[i2][j][p][q]
-15 * input[j2][h+p][w+q];}}
-16 store_output(voutput, output, i0);}
-Listing 4: Code of CNN found by Sisyphus
+To overcome these limitations, we introduce Sisyphus, a solution tailored for affine programs, where loop bounds and conditionals are affine functions of surrounding loop iterators and parameters (commonly referred to as Static Control Parts in the Polyhedral Model) (Girbal et al., 2006; Feautrier, 1988).
+Our objective is to define a search space that allows for the selection of various parameters, such as pragma insertion, loop ordering and tiling, within a single optimization problem.
+This approach avoids the need for backtracking or handling disjoint optimization problems.
+Our framework efficiently and rapidly explores this large design space, overcoming the limitations of other frameworks that either restrict the search space too much or explore a broad space with a DSE that only partially covers it.
+Most importantly, it guarantees that only valid code transformations are considered.
+To accomplish this, we propose a two-step process: first, the creation of a space tailored to each kernel, followed by the exploration of that space.
+The first step involves creating a template with fully distributed code that includes three levels of loops, each targeting specific optimizations.
+To achieve this, we strip-mined each original loop twice, ensuring that one of these loops is included at each level (assuming the permutation is valid).
+As a result, each original loop can be utilized across all three levels.
+For instance, in Listing 1, the loop $k$ (L2) is used in level 0 alongside loop L5 and in level 2 with loop L11 in Listing LABEL:lst:small_mem.
+Specifically, we have an innermost level for fine-grained unrolling, a level dedicated to pipelining, and another level focused on on-chip data caching based on available resources.
+As demonstrated in Listings LABEL:lst:small_mem and 2, depending on the resource constraints specified by the user, we obtain two different code versions.
+For instance, a reduced on-chip memory requirement necessitates the partial transfer of arrays $A$ and $B$ (line 17, 18 in Listing LABEL:lst:small_mem), leading to the use of a smaller on-chip buffer size that meets the user’s constraints.
+Once this template is generated, it establishes a search space where the unknowns consist of the loop bounds, loop ordering, buffer sizes for on-chip arrays, and the locations in the code where data transfers take place.
+Although other DSE methods can explore this space, we chose to use a cost model formulated as a Nonlinear Programming (NLP) problem.
+This approach allows us to efficiently navigate the theoretical space within seconds or minutes, leveraging accurate modeling made possible by compile-time analysis for affine programs.
+It is important to emphasize that exhaustive exploration using a simple cost model or synthesizing each design with HLS, as done in AutoDSE, would be impractical due to the sheer size of the search space.
+In contrast, the NLP approach enables us to explore this space quickly and efficiently, allowing for a more comprehensive investigation than traditional methods.
+For gemm, Sisyphus significantly enhances parallelism by splitting and permuting loops without relying on the reduction loop.
+In this configuration, the NLP identifies the setup shown in Listing 2.
+Both pipelined loops achieve an initiation interval (II) of 1, resulting in a throughput of 210 GFLOPs per seconds (GF/s).
+The reduction loop (L11) is partially unrolled, allowing the multiplications to be executed in parallel while performing four additions sequentially.
+This small unrolling, selected by the NLP, deepens the pipeline, improving the quality of results (QoR) even though the reduction must remain sequential.
+For the CNN, the NLP selects all parameters within the design space, including loop order after exploring 6!
+possibilities, data transfer points, on-chip array size, and the level of parallelism.
+This comprehensive analysis takes just 13 minutes and yields a throughput of 341 GF/s.
+Thanks to a well-defined space that only considers legal transformations and exploration guided by a cost model formulated as an NLP, our framework is able to select the theoretical optimum in the entire space in a matter of seconds or minutes and generate a design.
 3. Unified Space
@@
 This involves implementing maximal distribution for code transformation, followed by the design and implementation of a template capable of executing code transformation, tiles size selection and pragma insertion simultaneously.
-We illustrate our code transformation process step by step using the example provided in Listing 1.
-Maximum Distribution  To kickstart the process, we begin by thoroughly distributing the program, with the goal of maximizing the distance between each statement.
-This distribution strategy creates ample opportunities for parallelization by maximizing the separation between dependencies.
+We illustrate our code transformation process step by step using the example from Listings 1 to 2.
+Input  Sisyphus takes an affine C/C++ file as input and automatically extracts a polyhedral representation of the code using PoCC (PoCC, [n.
+d.]).
+This representation allows us to gather all relevant information from the code, such as loop order, loop bounds, array sizes, and more.
+We will use this information to create the template and to automatically generate the NLP discussed in Section 4.
+Our framework will be particularly effective for codes that are fully distributable and fully permutable.
+However, in cases where the transformations described later are not legally possible, our framework still functions but restricts the solution space more.
+Maximal Distribution  To kickstart the process, we begin by thoroughly distributing the program, with the goal of maximizing the distance between each statement.
+This distribution strategy creates ample opportunities for parallelization by maximizing the distance between dependencies (reverse of fusion).
 Typically, this results in one statement per loop body with perfectly nested loops.
@@
 It allows us to explore various schedule distributions and validate the legality of transformations by ensuring dependency constraints are preserved.
-After distribution, the example (Listing 1) yields two loop bodies.
-The first statement iterates through loops $i$ and $j$, while the second statement iterates through loops $i$, $k$, and $j$.
-Strip-mining  Following maximal distribution, we perform strip-mining on each loop twice.
-This process replaces the original loop, with a trip count $TC\_I0$, with three new loops, each having trip counts $TC\_I0\_0,TC\_I0\_1$ and $TC\_I0\_2$ with $TC\_I0\_0\times TC\_I0\_1\times TC\_I0\_2=TC\_I0$.
+The code presented in Listing LABEL:lst:distributed displays the fully distributed version of the code from Listing 1.
+Strip-mining  After maximal distribution, we apply strip-mining to each loop twice.
+This process transforms the original loop, such as the loop L0 in Listing LABEL:lst:distributed, into three loops with trip counts $I0\_S0$, $I1\_S0$, and $I2\_S0$ (L0, L1, L2 in Listing LABEL:lst:strip).
+The product of these three loop trip counts equals the original trip count, i.e., $I0\_S0\times I1\_S0\times I2\_S0=200$.
 Since strip mining is inherently legal, there is no necessity to validate the legality of this transformation.
 Subsequently, we can arrange the loops in different permutations if it is legal.
-Following the implementation of three levels of strip-mining, the first loop body of the code from Listing 1 undergoes transformation into the code depicted in Listing 5.
-1for (i0 = 0; i0 < TC_I0_0; i0++)
-2 for (i1 = 0; i1 < TC_I0_1; i1++)
-3 for (i2 = 0; i2 < TC_I0_2; i2++)
-4 for (j0 = 0; j0 < TC_J0_0; j0++)
-5 for (j1 = 0; j1 < TC_J0_1; j1++)
-6 for (j2 = 0; j2 < TC_J0_2; j2++)
-7 i = i0 * TC_I0_1 * TC_I0_2 + i1 * TC_I0_2 + i2;
-8 ...
-9 C[i][j] *= beta;
-Listing 5: Code of the first loop body of gemm with three level strip mining
-Following the explanation provided in the next paragraph, strip mining allows us to create opportunities for fine-grained unrolling, pipelining, sequential execution, coarse-grained unrolling and tiling if loop permutation is legal.
-This expands the range of possibilities, allowing each loop to integrate diverse hardware directives and be reorganized across different levels of the code.
-Loop Permutation  Next, we consider possible permutations of the strip-mined loops.
+After applying three levels of strip-mining, the code from Listing LABEL:lst:distributed is transformed into the code seen in Listing LABEL:lst:strip.
+Creation of different levels  Next, we consider possible permutations of the strip-mined loops.
 If these permutations are legal, we establish three loop levels.
 We ensure legality by checking dependency preservation using ISCC (Verdoolaege, 2011).
-The innermost level facilitates fine-grained unrolling (complete unrolling), which increases parallelism by duplicating statements and can utilize tree reduction if a reduction loop is unrolled and the option is enabled.
-The middle level facilitates pipelining, enhancing throughput by overlapping loop iterations.
-Finally, the outermost level coordinates coarse-grained unrolling, tiling, and/or sequential execution, optimizing to increase the parallelism, and controlling the size of the on-chip buffer through tiling.
-If the loop body consists of only one statement and loop permutation is feasible, we disable coarse-grained unrolling.
-This will achieve equivalent results because our template permits unrolling the same loop innermost for fine-grained unrolling.
-Furthermore, coarse-grained unrolling entails duplicating each array, including read-only arrays, to ensure full overlapping of the modules, even when there are no dependencies between the concurrently executed modules.
-In scenarios where permutations enable us to achieve perfectly nested loops capable of being pipelined, only one loop can usually be pipelined.
-Trying to pipeline all loops simultaneously, akin to pragma loop_pipelined, significantly complicates the design and extends synthesis times unreasonably.
-Thus, if a loop in the middle level is pipelined, it implies that the trip counts of the other loops (in the middle level) are reduced to one, ensuring manageable design complexity and synthesis times.
-As explained earlier, loop order selection is not necessary at the pipeline level.
-In the fine-grained level, arranging loops is straightforward; placing the reduction loop innermost enables efficient tree reduction.
-Conversely, at the coarse-grained or tiled level, prioritizing memory transfer latency is key.
-Therefore, we opt for a loop order that maintains the output stationary, minimizing costs associated with loading on-chip and storing off-chip during tile iterations, especially when the output is not fully transferred to the chip.
-In the context of our example, these transformations yield the code presented in Listing 3, with each level corresponding to specific transformations.
-It is notable that our framework maintains the original loop order.
-In Listing 3, if the loop on line 12 ($i2$) iterates once and the loop on line 3 ($i0$) iterates 200 times, we keep the original loop order, then proceed to pipeline and partially unroll loop $j$.
+Level 2:  The innermost level facilitates fine-grained unrolling (complete unrolling), which increases parallelism by duplicating statements and can utilize a tree reduction if a reduction loop is unrolled and the option is enabled.
+For example, this is illustrated by loops L9, L10, and L11 in Listing 2.
+Level 1:  The middle level facilitates pipelining, enhancing throughput by overlapping loop iterations.
+In cases where loop permutations allow us to achieve perfectly nested loops suitable for pipelining, we restrict pipelining to a single loop to ensure efficient synthesis.
+Flattening the loops at Level 1 into a single loop and then pipelining the resulting loop introduces significant design complexity and leads to excessively long synthesis times.
+Thus, if a loop at Level 1 is pipelined, it implies that the trip counts of the other loops at the same level are reduced to one, ensuring manageable design complexity and synthesis times.
+For instance, in Listing 2, the loop $j$ (L8) is pipelined for the second statement (S1).
+This results in the loops $k$ and $i$ (L10 and L11 in Listing 3 ) having a trip count of 1, enabling them to be eliminated from the code.
+Level 0:  Finally, the outermost level manages coarse-grained unrolling, tiling, and/or sequential execution.
+It not only controls parallelism at a coarse level but also determines the on-chip buffer size through tiling.
+As described in the following paragraph, the loop order at this level is included in the design space.
+This level is illustrated by loops L5, L6, and L7 in Listing LABEL:lst:small_mem.
+The loop bounds determine the on-chip buffer sizes for arrays A and B, transferred below loop L5.
+Resource constraints also influence the sequential execution of loop L7.
+If the loop body contains a single statement and loop permutation is feasible, we disable coarse-grained unrolling.
+This achieves equivalent results since our template allows fine-grained unrolling of the innermost loop.
+Loop Permutation  Loop order selection is not required for levels 1 and 2, as only a single loop is pipelined, and the innermost level (level 2) is fully unrolled, making the loop order irrelevant.
+Conversely, at the outermost level (level 0), all permutations are considered, and the NLP selects the permutation.
+In Listings LABEL:lst:small_mem and 2, we can see that the NLP selected two different permutations for the statement S1 (loops L5, L6, and L7).
+1int perm_S0[2] // Permutation order (0 for i0, 1 for j0)
+2int perm_S1[3] //Array to store permutation order (i0,j0,k0)
+3int lb_S0[2] = {I0_S0, J0_S0}; // loop bounds for S0
+4int lb_S1[3] = {I0_S1, J0_S1, K0_S1}; // loop bounds for S1
+5// possibilities to cache on-chip A, B and C
+6/********* Level 0 of S0 *********/
+7for (L0_S0=0; L0_S0 < lb_S0[perm_S0[0]]; L0_S0++) // L0
+8// possibilities to cache on-chip C
+9 for (L1_S0=0; L1_S0 < lb_S0[perm_S0[1]]; L1_S0++) // L1
+10 // possibilities to cache on-chip C
+11/********* Level 1 of S0 *********/
+12 for (j1=0; j1 < J1_S0; j1++) // L2
+13 for (i1=0; i1 < I1_S0; i1++) // L3
+14#pragma HLS pipeline
+15/********* Level 2 of S0 *********/
+16 for (i2=0; i2 < I2_S0; i2++) // L4
+17#pragma HLS unroll
+18 for (j2=0; j2 < J2_S0; j2++) // L5
+19#pragma HLS unroll
+20 C[i][j]*=beta; // S0
+21/********* Level 0 of S1 *********/
+22for (L0_S1=0; L0_S1 < lb_S1[perm_S1[0]]; L0_S1++) // L6
+23// possibilities to cache on-chip A, B and C
+24 for (L1_S1=0; L1_S1 < lb_S1[perm_S1[1]]; L1_S1++) // L7
+25 // possibilities to cache on-chip A, B and C
+26 for (L2_S1=0; L2_S1 < lb_S1[perm_S1[2]]; L2_S1++) // L8
+27 // possibilities to cache on-chip A, B and C
+28/********* Level 1 of S1 *********/
+29 for (j1=0; j1 < J1_S1; j1++) // L9
+30 for (k1=0; k1 < K1_S1; k1++) // L10
+31 for (i1=0; i1 < I1_S1; i1++) // L11
+32#pragma HLS pipeline
+33/********* Level 2 of S1 *********/
+34 for (i2=0; i2 < I2_S1; i2++) // L12
+35#pragma HLS unroll
+36 for (j2=0; j2 < J2_S1; j2++) // L13
+37#pragma HLS unroll
+38 for (k2=0; k2 < K2_S1; k2++) // L14
+39#pragma HLS unroll
+40 i = i0 * I1_S1 * I2_S1 + i1 * I2_S1 + i2;
+41 j = j0 * J1_S1 * J2_S1 + j1 * J2_S1 + j2;
+42 k = k0 * K1_S1 * K2_S1 + k1 * K2_S1 + k2;
+[…]
+[…]
+[…]
+[…]
+Listing 3: Template
 Fusion  We have opted not to incorporate fusion into our model for several reasons.
@@
 Therefore, we are willing to incur a minor cost, even if it involves transferring a tiled array multiple times.
-Space  Therefore, the challenge involves determining the trip counts of each loop (e.g., $TC\_I0\_0$).
+Template  In the context of our example, these transformations yield the code presented in Listing 3, with each level corresponding to specific transformations.
+The iterator of the loop retains its original name, with an added number indicating the level.
+For example, loop L14 in Listing 3 refers to loop L2 in Listing 1 at level 2 of the strip mining.
+Notably, our framework is capable of preserving the original loop order, allowing it to retain the existing loop structure if the code is already optimized.
+Space  Therefore, the challenge involves determining the trip counts of each loop (e.g., $J0\_S0$) and the loop order of the level 0.
 Additionally, the design must adhere to resource constraints, including the usage of DSP and on-chip memory.
 Hence, the trip counts corresponding to fully unrolled loops must be constrained to avoid over-utilization of DSP resources.
-The on-chip buffers must not exceed the on-chip memory capacity.
+The on-chip buffers size must not exceed the on-chip memory capacity.
 Consequently, for a large problem size, the outermost level will enable control over the buffer sizes.
-Thus, the space we consider encompasses the original schedule on which we can insert pragmas, but it also includes all the transformations mentioned previously for which the pragmas are inserted.
 4. NLP
-Presented here is a comprehensive set of constraints and variables employed in a non-linear program aimed at discovering the theoretical solution space outlined in Section 3.
+Presented here is a comprehensive set of constraints and variables employed in a nonlinear program aimed at discovering the theoretical solution space outlined in Section 3.
 We employ the methodology proposed by (Pouget et al., 2024b, a), adapted to our specific context.
@@
 This flexibility ensures adaptability across various platforms and compilers.
-To gather all the required information for the NLP, we utilize PoCC (PoCC, [n. d.]) to extract the intermediate representation.
+To gather all the required information for the NLP, we utilize PoCC (PoCC, [n.
+d.]) to extract the intermediate representation.
 4.1. Variables
-Tables 1 and 2 delineate the sets, variables, and constants utilized in our NLP formulation.
+Table 2 delineate the sets, variables, and constants utilized in our NLP formulation.
+4.2. Constraints
+Now, we explore the precise meaning and implications of each constraint.
+To make this more understandable, we provide examples using the listings presented in the paper.
+Trip Count  Equation 1 constrains the trip count of each loop, ensuring that the product of the trip counts equals the original trip count.
+The three loops after strip-mining, e.g., L6, L7, and L8 in Listing LABEL:lst:strip, must be equivalent to the original loop L2 in our example from Listing LABEL:lst:distributed.
+Therefore, we require $I0\_S0\times I1\_S0\times I2\_S0=200$.
+(1) $\forall l\in\mathcal{L},\prod_{v\in\mathcal{V}}TC_{l,v}=TC_{l}$
+Pipeline  Constraints 2, 3, 4, and 5 facilitate the selection of loop pipelining.
+Constraint 2 ensures that each loop has a boolean flag indicating whether the original loop is pipelined.
+For example, in Listings LABEL:lst:small_mem and 2, the statement $S1$ shows that the loop $j1$ (L8) is pipelined, while the original loop is $j$ (L3 in Listing 1), indicating that $loop_{j}\_pip=1$.
+As explained in Section 3, only one loop can be pipelined at each middle level.
+This means the sum of the boolean flags must be less than or equal to 1 (Eq.
+4), and the trip count of all other non-pipelined loops at the middle level must be 1 (Eq.
+3).
+For example, in the template (Listing 3), only loop L9 is pipelined at level 1 of S1 (loops L9, L10, and L11), so $loop_{j}\_pip=1$.
+Consequently, $loop_{i}\_pip=0$ and $loop_{k}\_pip=0$ (Eq.
+4), and Eq.
+3 indicates that the trip counts of loops L10 and L11 are 1, allowing their removal.
 Set Description
@@
 $\mathcal{L}_{s}$ the set of nested loops which iterate the statement $s$
-$\mathcal{L}_{s}^{red}$ the set of reduction loops which iterate the statement $s$
-$\mathcal{C}_{a_{d}}$ the set of loops which iterates the array $a$ at the dimension $d$
+$\mathcal{L}_{s}^{red}$ the set of reduction loops iterating the statement $s$
+$\mathcal{C}_{a_{d}}$ the set of loops iterating the array $a$ at dimension $d$
 $\mathcal{V}$ Level of the strip mining, 0 for coarse-grained/sequential, 1 for pipeline and 2 for fine-grained unrolled
@@
 $IL_{par},IL_{red}$ Iteration Latency of the operations without ($par$) and with ($red$) dependencies of the statement $s$
-$DSP_{s_{op}}$ Number of DSP used for the statement $s$ for the operation $op$
+$DSP_{s_{op}}$ Number of DSP used for the statement $s$ for the operation $op$ (This number accounts for the number of times the statement is replicated due to unrolling)
 $DSP_{available}$ Number of DSP available for the FPGA used
@@
 $reuse_{opt}$ Boolean for optimistic reuse
-Table 1. Overview of the set and constant employed in formulating the NLP
 Variable Description
@@
 $cache_{l}\_arr_{a}$ Boolean to know if the array $a$ is transferred on-chip after the loop $l$ at level 0 of the strip-mining
-Table 2. Overview of the variable employed in formulating the NLP
-4.2. Constraints
-Now, we explore the precise meaning and implications of each constraint.
-Trip Count  Equation 1 constrain the trip count of each loop, ensuring that the product of the trip counts equals the original trip count.
-(1) $\forall l\in\mathcal{L},\prod_{v\in\mathcal{V}}TC_{l,v}=TC_{l}$
+$perm_{s}$ ID of the permutation of the statement $s$
+$burst_{a}$ The maximum bitwidth at which the array $a$ can be transferred from off-chip to on-chip
+Table 2. Overview of the set, constant and variables employed in formulating the NLP
+Eq.
+5 allows the computation of the initiation interval (II), which is determined based on the loop’s properties and the iteration latency of the reduction operation.
+Specifically, when the loop being pipelined is a reduction loop, this follows the approach outlined in (Pouget et al., 2024b, a).
+In Listings LABEL:lst:small_mem and 2, the pipelined loops are non-reduction loops with no dependencies when iterating, hence $II=1$.
+However, if loop L10 in Listing 3 were pipelined, it would involve a reduction, meaning that at each iteration, C[i][j] needs to be read and written.
+Thus, to read for iteration $k+1$, we must wait for the write to finish at iteration $k$, which takes $IL_{+}$ cycles (latency of the reduction operation).
+(2) $\forall l\in\mathcal{L},loop_{l}\_pip\in\{0,1\}$
+(3) $\forall l\in\mathcal{L},(1-loop_{l}\_pip)\times TC_{l,1}==1$
+(4) $\forall s\in\mathcal{S},\sum_{l\in\mathcal{L}_{s}}loop_{l}\_pip\leq 1$
+(5) $\forall s\in\mathcal{S},II_{s}=\sum_{l\in\mathcal{L}_{s}}loop_{l}\_pip\times II_{l}$
 Coarse-grained unrolling  We enforce coarse-grained unrolling solely for non-reduction loops (Eq.
-4), where the unroll factor (UF) must divide and be less than or equal to the trip count of the current loop (Eq.
-5 and 2).
-As elaborated in Section 3, coarse-grained unrolling is disregarded if there is only one statement within the loop (Eq.
-3).
+8).
+For example, the loop L3 in Listing LABEL:lst:distributed cannot be coarse-grained unrolled as it is a reduction loop.
+Additionally, the unroll factor (UF) must divide the trip count of the current loop and be less than or equal to it (Eq.
+9 and 6).
+However, as elaborated in Section 3, coarse-grained unrolling is disregarded if there is only one statement within the loop (Eq.
+7).
+This implies that for Listing 1 coarse-grained unrolling is disabled.
 (6) $\forall l\in\mathcal{L},loop_{l}\_UF\leq TC_{l,0}$
@@
 (9) $\forall l\in\mathcal{L},loop_{l}\_UF\%TC_{l,0}==0$
-Pipeline  Constraints 6, 7, 8, and 9 facilitate the selection of loop pipelining, allowing only one pipeline per nested loop and computing the initiation interval (II) in accordance with the chosen pipelined loop.
-The II is determined based on the loop’s properties and the iteration latency of the reduction operation, specifically when the loop being pipelined is a reduction loop, following the approach outlined in (Pouget et al., 2024b, a).
-(6) $\forall l\in\mathcal{L},loop_{l}\_pip\in\{0,1\}$
-(7) $\forall l\in\mathcal{L},(1-loop_{l}\_pip)\times TC_{l,1}==1$
-(8) $\forall s\in\mathcal{S},\sum_{l\in\mathcal{L}_{s}}loop_{l}\_pip\leq 1$
-(9) $\forall s\in\mathcal{S},II_{s}=\sum_{l\in\mathcal{L}_{s}}loop_{l}\_pip\times II_{l}$
-On-chip memory  Constraints 10 and 11 ensure that each array is cached on-chip at a single location within the code, and that the on-chip tiled data fits within the available on-chip memory.
+On-chip memory  Constraints 10 and 11 ensure that each array, represented by a boolean (Eq.
+10), is cached on-chip at a single location within the same loop body, as only one boolean can be true for each loop body (Eq.
+11).
+In Listing 3, for instance, array $A$ can be cached either before any loops (line 5) or after L6, L7, or L8 (lines 23, 25, and 27).
+In Listing 2, $A$ is cached on-chip with the on-chip footprint matching the size of the original array (line 2).
+Meanwhile, in Listing LABEL:lst:small_mem, $A$ is cached on-chip at line 18, but with a smaller on-chip footprint compared to the original.
+Based on where each array is transferred, we can compute the on-chip footprint.
+For example, in Listing LABEL:lst:small_mem, array $A$ is transferred below loop L5, which has a trip count of 60, so the footprint of $A$ on-chip is $200\times\frac{240}{60}$.
 (10) $\forall l\in\mathcal{L},\forall a\in\mathcal{A},cache_{l}\_arr_{a}\in\{0,1\}$
+(11) $\forall a\in\mathcal{A},\forall s\in\mathcal{S},\sum_{l\in\mathcal{L}_{s}},cache_{l}\_arr_{a}=1$
 (12) $\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}cache_{l}\_arr_{a}\times ft\_array_{a}\_loop_{l}\leq Mem$
-Array partitioning  Equations 12, 13 and 14 guarantee that the maximum array partitioning is not exceeded, and it ensures that the partitioning of the array is greater than or equal to the maximum unroll factor applied to that dimension.
-Thanks to Equation 13, which ensures that each unroll factor ($TC_{l,2}$) divides $AP_{a,d}$, $AP_{a,d}$ directly determines the array partitioning applied in the generated code.
+Array partitioning  We consider unrolling only if all the data needed for unrolling can be accessed in parallel, i.e., these data are on different on-chip banks thanks to the pragma array_partitioning.
+For each array $a$ and each dimension $d$ of this array, we want the factor of the cyclic array partitioning for this dimension to be greater than the unroll factor (Eq.
+15) to ensure that all data are on different banks.
+If the array is reused, we need to ensure that the array partitioning is optimal for each statement.
+Therefore, we force the array partitioning to be a multiple of each unroll factor that iterates through this dimension (Eq.
+14).
+$AP_{a,d}$ gives the cyclic array partitioning factor for array $a$ in dimension $d$ that we use in the generated code.
+Eq.
+13 assures that the array partitioning per array (product of the partitioning factors per dimension) is less than or equal to the limit fixed by the user.
+For example, in Listing 2, $C$ is unrolled 200 times in the first dimension (lines 11, 12) and 4 times in the second dimension (lines 13, 14) for the first loop body, and 200 times in the first dimension (lines 25, 26) for the second loop body.
+Thus, we have $AP_{C,1}=200$ and $AP_{C,2}=4$, and $200\times 4=800\leq max_{part}$, where $max_{part}$ is 1,024 in our case.
 (13) $\forall a\in\mathcal{A},\prod_{d\in\mathbb{N}}AP_{a,d}\leq max_{part}$
@@
 (15) $\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}\geq TC_{l,2}$
-DSP utilization  Lastly, Equations 15, 16, 17, and 18 compute the DSP utilization under optimistic (full reuse) and pessimistic (no reuse) DSP reuse assumptions.
+DSP utilization  Finally, we need to manage DSP utilization.
+Given the challenges in predicting DSP reuse, we use a boolean flag, $reuse_{opt}$, set by the user to determine the level of DSP reuse.
+If $reuse_{opt}=1$, we assume an optimistic reuse scenario: DSPs used in one loop body can be directly reused by subsequent loop bodies if the operations are identical.
+However, if the operations differ, as discussed in (Pouget et al., 2024b, a), no reuse is assumed.
+Therefore, the total DSP usage is calculated as the sum of the maximum DSPs required for each operation across all loop bodies (Eq.
+16).
+If $reuse_{opt}=0$, we assume no reuse between different loop bodies, and we sum the DSP requirements for each statement (Eq.
+17).
+Based on the value of $reuse_{opt}$, the total DSP usage must be less than or equal to the available DSPs (Eq.
+18 and 19).
+For instance, if a multiplication operation (*) uses 3 DSPs and an addition operation ($+$) uses 2 DSPs, in Listing 2, S0 requires $200\times 4\times 3$ DSPs for multiplications, and S1 requires $200\times 4\times 3+200\times 4\times 3+200\times 4\times 2$ DSPs for two multiplications and one addition.
+With optimistic reuse for the same operation (e.g., multiplication), the DSP usage is calculated as: $\max(200\times 4\times 3,2\times 200\times 4\times 3)+200\times 4\times 2$.
+With pessimistic reuse, the DSP usage is: $200\times 4\times 3+2\times(200\times 4\times 3)+200\times 4\times 2$.
 (16) $DSPs\_used_{opt}=\sum_{op\in\{+,-,*,/\}}\max_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})$
@@
 We structure the objective function akin to the approach described in (Pouget et al., 2024b, a), tailoring it to meet the specific needs of our problem.
-Here, $Lat_{2}$ corresponds to the fine-grained unrolled tile, $Lat_{1}$ denotes the pipeline tile that incorporates the fine-grained unrolled tile, and recursively, $Lat_{0}$ encompasses $Lat_{1}$, enabling coarse-grained unrolling and facilitating on-chip data caching.
-$\left.\begin{cases}&Lat_{2}=IL_{par}+IL_{seq}\times\prod_{l\in\mathcal{L}^{red}}\times\log_{2}(TC_{l,2})\\ &Lat_{1}=Lat_{2}+II\times(TC_{l,1}-1)\\ &Lat_{0}=\prod_{l\in\mathcal{L}}\frac{TC_{l,0}}{loop_{l}\_UF}\times Lat_{1}\\ &L_{mem}=\sum_{l\in\mathcal{L}}\max_{a\in\mathcal{A}}(cache_{l}\_arr_{a}\times ft\_array_{a}\_loop_{l})\\ &obj\_func=Lat_{0}+L_{mem}\end{cases}\right.$
-5. Code Generation and Optimization
-The NLP allows us to find the trip count of each loop, the array partitioning for each array, and the size of the on-chip buffers.
-By considering the loop order and the pragmas applied to each loop as described in Section 3, we can directly generate the code.
-We automatically generate the functions that load data from off-chip to on-chip with the maximum burst size possible and vice versa with the functions that store the data off-chip.
-Additionally, we incorporate further optimizations to enhance the QoR.
-5.1. Optimization for non-constant trip count
-To optimize loops with non-constant trip counts, we implement a code transformation that preserves the NLP’s estimation while simplifying compilation for the HLS compiler.
+Here, $Lat^{s}_{2}$ corresponds to the fine-grained unrolled level (level 2) for the statement $s$, $Lat^{s}_{1}$ denotes the pipeline tile that incorporates the fine-grained unrolled level (level 1), and recursively, $Lat_{0}$ encompasses $Lat_{1}$, enabling coarse-grained unrolling and facilitating on-chip data caching (level 0).
+The memory latency $Lat^{s}_{mem}$ for a statement $s$ corresponds to the time required to transfer the array from off-chip to on-chip and back for the arrays needed by a statement $s$.
+We assume the memory transfer is pipelined, allowing the transfer of one data element (with a maximum bitwidth of 512 bits) per cycle.
+Therefore, the latency at a given level is calculated as the array footprint divided by the burst size (bitwidth used for transferring the array), multiplied by the trip count of the loop that iterates over the function responsible for the transfer, based on the selected permutation of level 0.
+Additionally, if an array is fully transferred on-chip for a statement $s$ and reused by a subsequent statement $s^{’}$, no further transfers are needed.
+The burst size for an array must be a divisor of its last dimension.
+For instance, if we have an array $float\text{ }A[512][20]$, the maximum burst size of this array $A$ is 128 bits, since the last dimension, $20\times 32$ bits, can be divided by the largest power of 2, which is 128.
+$\left.\begin{cases}&Lat^{s}_{2}=IL_{par}+IL_{seq}\times\prod_{l\in\mathcal{L}^{red}}\times\log_{2}(TC_{l,2})\\ &Lat^{s}_{1}=Lat^{s}_{2}+II\times(TC_{l,1}-1)\\ &Lat^{s}_{0}=\prod_{l\in\mathcal{L}}\frac{TC_{l,0}}{loop_{l}\_UF}\times Lat^{s}_{1}\\ &L^{s}_{mem}=\sum_{l\in\mathcal{L}}\max_{a\in\mathcal{A}}(cache_{l}\_arr_{a}\times ft\_array_{a}\_loop_{l}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\times TCs(perm_{s})/burst_{a})\\ &Lat_{s}=Lat^{s}_{0}+L^{s}_{mem}\end{cases}\right.$
+Finally, the objective function (Eq. 20) is defined as the sum of the latencies of each loop body.
+(20) $obj\_func=\sum_{s\in\mathcal{S}}Lat_{s}$
+5. Post-Optimization and Usage
+The NLP determines the trip count for each loop, array partitioning, and on-chip buffer sizes.
+Functions for loading/storing data between off-chip and on-chip memory are automatically generated with the maximum burst size.
+Further optimizations are also applied to improve QoR.
+Optimization for Non-Constant Trip Count  To optimize loops with non-constant trip counts, we implement a code transformation that preserves the NLP’s estimation while simplifying compilation for the HLS compiler.
 We achieve this by replacing loops with non-constant trip counts with the maximal trip count computed using PoCC (PoCC, [n.
 d.]).
-Subsequently, we create a function for all statements iterated by this loop, ensuring compliance with the constraints of the non-constant trip count loop.
-This function returns the computation if the constraints are met or returns just the value of the output otherwise.
-Introducing a condition solely above the statement results in excessive compilation times.
+Next, we replace all statements iterated by this loop with calls to a function.
+For example, instead of using $C[i][i]+=B[i][i]$, we replace it with $C[i][j]=f(C[i][j],B[i][i])$, where the function $f$ ensures compliance with the constraints of the non-constant trip count loop.
+This function returns the computed value if the constraints are satisfied; otherwise, it simply returns the output value.
 Therefore, we employ these techniques to reduce compilation time and achieve designs with a good QoR.
-5.2. Overlapping computation and communication
-The constraints of the NLP consider the sum of the footprints of all arrays used during the design execution.
-Thus, we have the option, while already anticipating resource constraints, to overlap the transfer of arrays between off-chip and on-chip (or vice versa) at the same level or between the transfer of the arrays needed for the next loop body with the computation of the current loop body.
-To achieve this, we separate each loop body into different functions, as well as the functions that transfer input data between off-chip and on-chip.
-Additionally, we create new functions that execute these independent functions in parallel.
-To ensure parallel execution, within these newly created functions designed for overlap, we include only independent functions.
-5.3. Code transformation for HLS
-To simplify the understanding of the HLS compiler, particularly to ensure that the compiler can pipeline loops with the correct initiation interval (II), we simplify the reductions (regardless of whether we use tree reductions) in the fine-grained unrolled part.
-As specified in section 3, the fully unrolled loops are placed innermost.
-Therefore, before the reduction, we create a variable to accumulate the reduction, which we then add to the output.
-Additionally, we incorporate the pragma loop_flatten off on all reduction loops above the pipeline.
-While theoretically, this adjustment could enhance the QoR, in practice, the compiler might choose not to pipeline at all if it cannot flatten the loop and pipeline correctly.
+Grouping Data Transfers  In the NLP, we already account for parallel transfers of arrays when they are transferred at the same level in the code (e.g., within the same loop).
+However, we do not consider parallel transfers between arrays belonging to different loop bodies.
+In the 2mm example, we first handle the transfer of arrays required for the initial matrix multiplication, only transferring the arrays for the second multiplication after the first is complete.
+However, FPGA DRAM banks enable parallel transfers of different arrays.
+Therefore, we can optimize load and store operations during post-processing to enhance memory transfer overlap.
+Code Transformation for HLS  To simplify the understanding of the HLS compiler, particularly to ensure that the compiler can pipeline loops with the correct initiation interval (II), we simplify the reductions (regardless of whether we use tree reductions) in the fine-grained unrolled part (level 2).
+Therefore, under the pipeline we create a variable to accumulate the reduction, which we then add to the output.
+Additionally, we use the loop_flatten off pragma on all reduction loops above the pipeline.
+Although this adjustment could theoretically improve the QoR, in practice, flattening these loops can complicate the code and potentially increase the initiation interval (II) of the pipeline.
+Usage  Our framework is designed to be explainable and user-friendly, requiring the user to provide only the affine C/C++ code of the kernel to be optimized and the available resources.
+To fully benefit developers, our NLP formulation allows the template to be completely customized to the user’s needs.
+Each variable in the design space can be defined by the user if they have specific constraints.
+For example, in the gemm template (Listing 3), the user can set the permutations for level 0 while allowing the NLP to explore other parameters.
+Additionally, they can specify the on-chip size of matrix A, enabling exploration within the defined constraints.
 6. Evaluation
+The goal of this evaluation is to demonstrate how effectively our approach applies code transformations and inserts hardware directives while respecting user-defined resource constraints.
+We compared our method against four other frameworks: AutoDSE (Sohrabizadeh et al., 2021), NLP-DSE (Pouget et al., 2024b), HARP (Sohrabizadeh et al., 2023), and ScaleHLS (Ye et al., 2022).
+Since generating a bitstream without significant manual intervention is challenging for some frameworks, we first demonstrate our ability to achieve a high-quality design by utilizing all available resources—ensuring fair comparison with the other frameworks—without considering place and route.
+We generate HLS reports using the vitis-flow, which accounts for memory transfers between off-chip and on-chip, necessitating the inclusion of memory transfer steps in ScaleHLS’s code.
+Thus, we show that our designs are synthesizable by targeting the resources of a single SLR, which eliminates the need for manual partitioning.
+Due to the time required for bitstream generation, we restrict our evaluation to a subset of the kernels.
 6.1. Setup
-We employ kernels sourced from Polybench/C 4.2.1 (PolyBench, [n.
-d.]), supplemented by a Convolutional Neural Network (CNN) kernel and matrix multiplication tasks akin to those in the BERT transformer model.
-Our computations utilize single-precision floating-point data types as the default, facilitating comparison with AutoDSE and NLP-DSE.
-We operate on medium-sized datasets from PolyBench/C (PolyBench, [n.
-d.]).
-We also compare our approach with HARP (Sohrabizadeh et al., 2023), which seeks to enhance High-Level Synthesis design exploration by integrating a Graph Neural Network model that predicts HLS tool behavior.
-They conduct exploration within this space utilizing the GNN model for one hour and then synthesize the top-10 designs.
+We conducted experiments using kernels from Polybench/C 4.2.1 (PolyBench, [n.
+d.]), a CNN kernel, and matrix multiplication tasks similar to those in the BERT transformer model.
+All computations used single-precision floating-point data for direct comparison with frameworks like AutoDSE, NLP-DSE, and ScaleHLS, using medium-sized Polybench/C datasets.
+We also compare our approach with HARP (Sohrabizadeh et al., 2023), which enhances HLS design exploration using a Graph Neural Network (GNN) model to predict HLS tool behavior.
+HARP explores this space for one hour and synthesizes the top 10 designs.
 We use the medium-sized Polybench kernels found in their training set with double-precision floating-point data types, aiming to deploy HARP under optimal conditions.
-As detailed in (Pouget et al., 2024b, a), HARP’s effectiveness is not universally applicable without the necessary fine-tuning or training, particularly when confronted with diverse problem sizes.
-We chose these problem sizes to demonstrate the efficacy of our approach in transforming codes and inserting pragmas in scenarios where fully unrolling is not feasible.
-Additionally, we aim to illustrate the effectiveness of our framework in tiling code for substantial tasks such as CNN and bert_3072_100.
-This approach ensures that only a portion (a tile) of each array fits on-chip, aligning with the available memory constraints.
-The problem size and original loop order of CNN are I,J=256, H,W=224 and P,Q=5.
-For bert_n_m matrix multiplication the schedule and problem size are I=n, J=m, K=n (reduction).
-We utilize the AMD/Xilinx Vitis HLS compiler (Xilinx, 0232) to showcase the effectiveness of Sisyphus.
-Report generation is performed using AMD/Xilinx Vitis 2023.2.
-When enabling tree reduction, we utilize the ”funsafe math optimizations” option to enable commutative/associative reduction operators, thereby facilitating reductions to be implemented in logarithmic time.
-In our comparison with HARP (Sohrabizadeh et al., 2023), we utilized the same parameters as the authors, employing Vitis 2021.1 and activating the option for tree reduction.
-As our designated hardware platform, we deploy the Xilinx Alveo U200 device with a target frequency of 250 MHz.
-We conduct an analysis of the kernels and automatically generate each NLP problem using PoCC (PoCC, [n.
-d.]).
-Solving these NLP problems, we employ the AMPL description language and utilize the commercial BARON solver version 21.1.13 (Tawarmalani and Sahinidis, 2005; Sahinidis, 2017).
-For our experimental setup, we harness the computing power of 2 AMD EPYC 7V13 64-Core Processor.
+As detailed in (Pouget et al., 2024b), HARP’s effectiveness is not universally applicable without the necessary fine-tuning or training, particularly when confronted with diverse problem sizes.
+We selected these problem sizes to highlight how our method performs in transforming code and inserting pragmas in cases where full unrolling is not feasible.
+We also demonstrate the effectiveness of our framework in tiling large tasks such as CNN and bert_3072_100, ensuring that only a small portion (a tile) of each array fits on-chip, in line with memory constraints.
+For the CNN kernel, the problem size and loop dimensions are the same as described in Section 2.
+For bert_n_m matrix multiplication, the dimensions are I = n, J = m, K = n (reduction).
+We used the AMD/Xilinx Vitis HLS compiler (Xilinx, 0232) to demonstrate our method’s effectiveness, generating reports with Vitis 2023.2.
+When enabling tree reduction, we selected the “funsafe math optimizations” option to allow commutative and associative reduction operators for logarithmic time reductions.
+For our hardware setup, we targeted the Xilinx Alveo U200 device running at 250 MHz.
+We use the commercial solver Gurobi 11.0.0 (Gurobi Optimization, LLC, 2024) with AMPL.
+The experiments were performed on an AMD EPYC 7V13 64-core processor, with 32 GB of memory and 4 cores allocated for each run.
 6.2. Experimental Evaluation
-The aim of this evaluation is to demonstrate our capability to perform the proper transformation while simultaneously inserting the correct hardware directives.
-To achieve this, we compare our approach with three frameworks, that conduct design space exploration to insert pragmas with a fixed schedule.
-The input codes provided are the original codes from Polybench or as described in the subsection 6.1.
-However, we modify the kernels containing loops with non-constant trip counts using the method outlined in Section 5.3.
-This adjustment ensures a fair evaluation with the transformations we apply.
-We compare our method with AutoDSE (Sohrabizadeh et al., 2021) and NLP-DSE (Pouget et al., 2024a, b) described in Section 2, and HARP (Sohrabizadeh et al., 2023).
-We generate the AutoDSE space automatically by running the command ds_generator.
-We replace the unroll factor and tile size with all factors of the trip count to ensure consistency within the space.
-However, since maximum array partitioning is not a parameter within AutoDSE’s space, this space, as observed later with the kernel atax, is slightly larger.
-AutoDSE is executed using the bottleneck method with a DSE timeout set to 1,000 minutes and a timeout of 180 minutes for each HLS synthesis.
-NLP-DSE is executed with the parameters specified in the paper, including a 30-minute timeout for the BARON solver and the space provided as input is $\{\infty,2048,1024,512,256,128,64,32,16,8,1\}$.
-For HARP, we adhere to the parameters outlined by the authors mentioned earlier.
-In evaluating Sisyphus, we set a solver timeout of 14,400 seconds (4 hours) for the Baron solver, which generates only one design.
-By default, we assume that the compiler can efficiently reuse resources between statements that are not executed in parallel.
-If the synthesis results in the design exceeding DSP utilization, we relaunch the NLP with a pessimistic DSP utilization assumption—meaning we consider no reuse between statements—and regenerate the design accordingly.
-For heat-3d, symm, syrk, syr2k, and bert_3072_100 we had to apply the constraint indicating no reuse (Eq.
-16), as the constraint with optimistic reuse (Eq.
-15) resulted in kernels with resource over-utilization.
-The maximum partitioning constraint, $max_{part}$ (cf.
-Section 4), is set at 1024, as determined by our use of AMD/Xilinx compilers and FPGAs, adhering to their specified limit.
-Each synthesis process begins with a C-simulation to verify the correctness of the generated code.
+To ensure a fair comparison, we modified kernels with non-constant trip counts for all benchmarks, as detailed in Section 5, except for ScaleHLS, which already handles such loops.
+We generated AutoDSE’s design space using the ds_generator command, applying all trip count factors for unroll and tile sizes.
+Since AutoDSE lacks array partitioning, its design space is slightly larger.
+AutoDSE was run with a 1,000-minute exploration timeout and a 180-minute HLS synthesis timeout.
+For NLP-DSE, we followed the original paper’s parameters.
+HARP was run with the authors’ settings using Vitis 2021.1 and tree reduction enabled.
+ScaleHLS assumes that all kernel data are already located in on-chip memory.
+To provide a fair comparison, we modified their code to handle off-chip to on-chip memory transfers.
+We optimized these transfers to 512 bits per cycle, with all arrays transferred in parallel.
+Each array was assigned to a different DRAM bank.
+However, due to the extensive manual modifications required, we did not extend this comparison to all kernels.
+For Sisyphus evaluation, we set the Gurobi solver timeout to 14,400 seconds (4 hours) per design.
+By default, we assume optimistic resource reuse between non-parallel statements.
+If DSP utilization exceeds limits, we rerun the NLP with pessimistic assumptions (no resource reuse) and regenerate the design.
+The maximum partitioning constraint, $max_{part}$, was set to 1024, per AMD/Xilinx guidelines.
+Each synthesis starts with a C-simulation to verify correctness.
+We generate Vitis reports using the vitis-flow, which counts memory transfers for all frameworks and considers the resources of the entire FPGA.
+For bitstream generation, we consider only one SLR to eliminate any manual intervention.
+We consider 60% of the resources of a single SLR to quickly identify a synthesizable design, minimizing the number of times AutoDSE needs to be run.
 6.3. Comparison of the Methods
-6.3.1. AutoDSE and NLP-DSE
-Table 3 provides a comparison among AutoDSE, NLP-DSE, and Sisyphus.
-The TR column denotes the status of the tree reduction option, while the T column indicates the specific transformations applied, with D representing distribution, T for data tiling, and P for permutation.
-Across all evaluated kernels, we consistently achieve comparable or superior performance, with the exception of the atax kernel, where AutoDSE yields designs approximately 1% to 2% faster, and heat-3d, where NLP-DSE makes a 1% improvement.
-However, the designs obtained by AutoDSE fall outside our defined space due to the stringent constraint of a maximum array partitioning limit of 1024 imposed by AMD/Xilinx Vitis.
-The design found by NLP-DSE for heat-3d shares the same schedule and pragma as ours, but slight enhancements in QoR are attributed to how Merlin handles memory transfer.
-{adjustwidth}
--1in-1in
-GF/s Perf. Improvement
-Kernel TR T Sisyphus AutoDSE NLP-DSE
-atax 1 D 1.96 0.99x 1.00x
-atax 0 D 1.86 0.98x 1.00x
-heat-3d - - 131.33 35.09x 0.99x
-jacobi-2d - - 128.83 13.06x 6.60x
-2mm 1 D,P 164.34 403.55x 1.33x
-2mm 0 D,P 162.24 775.26x 1.36x
-bert_100_64 1 D,P 88.30 1.08x 1.03x
-bert_100_64 0 D,P 76.76 1.09x 1.09x
-bert_100_768 1 D,P 218.08 1.04x 1.10x
-bert_100_768 0 D,P 216.95 1.13x 1.13x
-bert_100_3072 1 D,P 220.29 1.10x 1.10x
-bert_100_3072 0 D,P 219.97 1.11x 1.11x
-gemm 1 D,P 198.26 1.79x 1.51x
-gemm 0 D,P 210.63 9.25x 9.57x
-bicg 1 D,P 1.96 1.97x 1.97x
-bicg 0 D,P 1.86 1.92x 1.93x
-gemver 1 P 17.85 4.64x 1.77x
-gemver 0 P 12.45 30.91x 7.24x
-mvt 1 P 13.71 1.76x 1.76x
-mvt 0 P 10.00 1.45x 1.45x
-bert_3072_100 1 D,P,T 42.93 2.98x 3.31x
-bert_3072_100 0 D,P,T 44.67 536.42x 626.60x
-cnn 1 D,P,T 340.20 8.00x 8.91x
-cnn 0 D,P,T 339.34 8.05x 10.67x
-doitgen 1 D,P 53.22 1.33x 2.66x
-doitgen 0 D,P 54.05 5.46x 2.71x
-symm 1 D,P 236.82 8.15x 7.20x
-symm 0 D,P 236.56 8.14x 7.19x
-syr2k 1 D,P 428.12 9.28x 3.12x
-syr2k 0 D,P 428.39 17.57x 3.13x
-syrk 1 D,P 306.07 12.44x 4.40x
-syrk 0 D,P 308.71 12.55x 4.43x
-Average 153.65 59.99x 22.85x
-Geo Mean 69.09 5.97x 2.89x
-Table 3. Comparison of the Throughput (GF/s) achieved with Sisyphus, AutoDSE and NLP-DSE
-Sisyphus averages a speedup of 59.99x and 22.85x over AutoDSE and NLP-DSE, respectively, across the evaluated kernels. In terms of geometric mean, Sisyphus achieves a speedup of 5.97x and 2.89x.
-Conserve the original schedule when needed  In the case of atax and jacobi-2d, Sisyphus preserves the original schedule and solely inserts pragmas.
-Thus, it effectively maintains the existing schedule while efficiently incorporating pragmas, a process comparable to that of two DSEs dedicated to pragma insertion.
-Regarding heat-3d, the original design opts for partial unrolling of the three loops with unroll factors of 2, 2, and 38, respectively, enabling a level of parallelism unattainable without code transformation.
-However, upon evaluation, we found that this design was over-utilizing resources.
-Therefore, the NLP solution, using a pessimistic constraint for DSP reuse, resulted in a design with the original schedule.
-Although we note improvements for jacobi-2d, it is noteworthy that the same pragma insertion was utilized in at least one DSE.
-However, despite this consistency, AMD/Xilinx Vitis 2023.2 failed to synthesize the design generated by the source-to-source compiler Merlin, which the two DSEs employ.
-Code transformation to manage reduction loop  In the cases of 2mm, bert, and gemm, we observe notable improvements resulting from code transformations, particularly in effectively managing the reduction loop.
-This enhancement is particularly evident with bert_3072_100, where the reduction loop exhibits a large trip count, as well as in gemm when tree-reduction is not used.
-Code transformation with tiling  The selection of tile size becomes crucial for determining which array sizes, particularly for bert_3072_100 and CNN with their large arrays, should be cached on-chip and where they should be transferred.
-Organizing the outermost loop as output stationary (cf.
-Section 5.3), prompts the NLP to perform partial transfers of the array iterated by these loops in these examples.
-Consequently, the NLP’s choices in these scenarios aim to optimize data reuse and minimize redundant memory transfers.
-However, for CNN, its large size requires multiple transfers of the array input (cf.
-Listing 4).
+6.3.1. AutoDSE, NLP-DSE and Scale-HLS with tree reduction
+Table 3 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE), Scale-HLS (S-HLS) and Sisyphus (ours) with tree reduction.
+The T column indicates the specific transformations applied, with D representing distribution, T for data tiling, and P for permutation.
+The column Perf.
+Impr.
+of ours framework vs.
+displays the performance improvement of our framework compared to the three different frameworks.
+Across all evaluated kernels, we consistently achieve performance that is comparable to or better than alternatives, except for the atax, and gesummv kernels, where we observe a slowdown of less than 5%.
+For 3mm, symm and trmm we had to apply the constraint indicating no reuse (Eq.
+17), as the constraint with optimistic reuse (Eq.
+16) resulted in kernels with resource over-utilization.
+For Gesummv, ScaleHLS is 1.41x faster.
+However, 91% of our design’s execution time is spent on data transfers.
+ScaleHLS benefits from 512-bit data transfers, which we manually applied for testing.
+Our framework uses 64-bit chunks, the largest bitwidth that divides the array.
+With 512-bit transfers, our framework could achieve 14.50 GF/s.
+For other memory-bound kernels like Atax and Gesummv, the slight performance differences observed with other frameworks are largely due to how Merlin handles memory transfers.
+GF/s Perf. Impr. of ours framework vs.
+Kernel T Ours A-DSE N-DSE S-HLS
+2mm D,P 175.49 431.17x 1.42x 5.50x
+3mm D,P 140.83 80.75x 1.02x 3.86x
+atax D 1.96 0.99x 1.00x 1.23x
+bicg D,P 1.96 1.97x 1.97x 1.18x
+doitgen D,P 54.05 1.35x 2.71x 8.45x
+gemm D,P 203.39 1.84x 1.55x 5.81x
+gemver P 17.62 4.58x 1.74x 9.39x
+gesummv D 1.96 0.99x 1.00x 0.71x
+mvt P 13.24 1.70x 1.70x 1.26x
+symm D,P 240.50 8.28x 8.28x 2,704.51x
+syr2k D,P 254.69 5.52x 1.86x 174.47x
+syrk D,P 161.20 6.55x 2.32x 358.95x
+trmm D,P 148.53 8,345.10x 8,312.42x 2,376.91x
+Average 108.88 683.91x 641.46x 434.78x
+Geo Mean 40.13 9.17x 3.47x 16.15x
+Table 3. Throughput (GF/s) Comparison of Sisyphus, AutoDSE, NLP-DSE, and ScaleHLS with Tree Reduction Using Vitis HLS
+Sisyphus averages a speedup of 683.91x, 641.46x and 434.78x over AutoDSE, NLP-DSE and Scale-HLS respectively, across the evaluated kernels.
+In terms of geometric mean, Sisyphus achieves a speedup of 9.17x, 3.47x and 16.15x.
+Preserve the original schedule when necessary  For atax, Sisyphus preserves the original schedule and inserts pragmas, effectively maintaining the existing schedule while efficiently incorporating pragmas, similar to two DSEs focused on pragma insertion.
 Code transformation for memory-bound kernel  For mvt and gemver, both memory-bound kernels, effective optimization entails loop permutations to prevent redundant array transfers and accommodate array partitioning within the constraints of the AMD/Xilinx compiler’s 1024 limit.
 Sisyphus adeptly manages these loop permutations and pragma insertions, guaranteeing optimized performance for these kernels.
-Additionally, in the case of gemver without tree reduction, the NLP reverses all loop orders to pipeline the non-reduction loop, further enhancing the QoR.
 For bicg, fully distributing and arranging the loops appropriately can enhance parallelization capabilities.
-Allow to achieve a parallelism not achievable without loop transformation  For doitgen, symm, syrk, and syr2k, the original schedule imposes constraints on the achievable level of parallelism.
-This limitation elucidates the performance enhancements facilitated by Sisyphus.
-In the case of syrk and syr2k, the primary advantageous transformation is the maximal loop distribution.
-Meanwhile, for doitgen and symm, we observe QoR improvements because code transformations enable a combination of loop unrolling not attainable even with maximal loop distribution.
+Allow to achieve a parallelism not achievable without loop transformation  For doitgen, symm, syrk, syr2k, and trmm, the original schedule imposes constraints on the achievable level of parallelism.
+This limitation highlights the performance enhancements made possible by Sisyphus.
+We observe QoR improvements for these kernels because the code transformations enable a combination of loop unrolling that cannot be achieved even with maximal loop distribution, as each loop can be partially unrolled.
+6.3.2. AutoDSE and NLP-DSE without tree reduction
+Transformations without tree reduction, which may be needed to preserve floating-point data precision, must be different since the optimizations differ when tree reductions are not involved.
+Table 4 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE) and Sisyphus (ours) without tree reduction.
+Sisyphus achieves an average speedup of 676.03x over AutoDSE and 567.53x over NLP-DSE across the evaluated kernels without tree reduction.
+The geometric mean speedup is 18.50x and 4.80x, respectively.
+On average, the performance improvement without tree reduction is greater than with tree reduction, particularly for gemm (Listing 2) and gemver.
+This is because, without tree reduction, optimizing the reduction loop becomes challenging, making code transformations even more necessary.
+However the difference observed between syrk and syr2k with tree reduction is due to the HLS compiler achieving an II=2 instead of the expected II=1.
+However, without tree reduction, the compiler successfully achieves II=1.
+GF/s Perf. Impr. Ours vs.
+Kernel T Ours A-DSE N-DSE
+2mm D,P 171.07 817.42x 1.44x
+3mm D,P 140.11 511.64x 1.04x
+atax D 1.86 0.98x 1.00x
+bicg D,P 1.86 1.92x 1.93x
+doitgen D,P 54.05 5.46x 2.71x
+gemm D,P 210.61 9.25x 9.57x
+gemver P 15.01 37.28x 8.73x
+gesummv D,P 1.81 0.97x 0.97x
+mvt P 9.75 1.42x 1.42x
+symm D,P 240.55 8.28x 7.32x
+syr2k D,P 429.34 17.61x 3.14x
+syrk D,P 333.03 13.54x 4.78x
+trmm D,P 130.90 7,362.68x 7,333.85x
+Average 133.84 676.03x 567.53x
+Geo Mean 41.62 18.50x 4.80x
+Table 4. Throughput (GF/s) Comparison of Sisyphus, AutoDSE and NLP-DSE without Tree Reduction Using Vitis HLS
+6.3.3. AutoDSE and NLP-DSE for Bert and CNN
+Table 5 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE) and Sisyphus (ours).
+The TR column denotes the status of the tree reduction option.
+For bert_100_64, we observed a minor decrease in QoR, attributed to the NLP selecting a configuration where the pipeline was not applied as expected.
+Code transformation with tiling  The selection of loop order and tile size is critical for determining which array sizes, particularly for large problem size like bert_3072_100 and CNN layers, should be cached on-chip and where they should be transferred.
+In these instances, the NLP aims to enhance data reuse and minimize unnecessary memory transfers while identifying the theoretical optimal balance with parallelism.
+GF/s Perf. Impr. Ours vs.
+Kernel TR T Ours A-DSE N-DSE
+bert_100_64 1 D,P 81.22 0.99x 0.95x
+bert_100_64 0 D,P 76.76 1.09x 1.09x
+bert_100_768 1 D,P 218.08 1.12x 1.19x
+bert_100_768 0 D,P 216.95 1.13x 1.13x
+bert_100_3072 1 D,P 241.59 1.21x 1.21x
+bert_100_3072 0 D,P 241.06 1.22x 1.22x
+bert_3072_100 1 D,P,T 314.89 21.85x 24.29x
+bert_3072_100 0 D,P,T 82.41 989.65x 1,156.02x
+cnn 1 D,P,T 314.15 7.38x 8.23x
+cnn 0 D,P,T 341.20 8.09x 10.73x
+Average 212.83 103.37x 120.61x
+Geo Mean 185.32 4.38x 4.69x
+Table 5. Throughput (GF/s) Comparison of Sisyphus, AutoDSE and NLP-DSE Using Vitis HLS
 6.3.4. HARP
@@
 Regarding mvt, HARP produced comparable performance to our own due to their use of the Merlin source-to-source compiler, which facilitates loop permutations when partially unrolling consecutive loops.
-This allows for them to find the same schedule that our framework identified.
+This allows them to find the same schedule that our framework identified.
 However, for Gemm, the design discovered by HARP falls outside our specified parameter space, particularly because of our restriction on maximal array partitioning.
-Nonetheless, it is worth noting that it is the only design among the top 10 evaluated that demonstrates a performance enhancement over our framework.
-GF/s Resource Utilization (%) Perf. Imp
+Notably, it is the only design in the top 10 that outperforms our framework.
+GF/s Resource Utilization (%) Perf. Impr.
 BRAM,DSP,LUT,FF
-Kernel HARP Sisyphus HARP Sisyphus
-atax 1.72 1.77 78,52,49,50 38,26,25,31 1.03x
-bicg 0.92 1.77 75,25,35,36 38,26,24,26 1.92x
-gemm 125.59 99.96 29,80,55,40 1,67,51,34 0.80x
-gemver 1.66 7.38 29,28,35,19 43,27,39,47 4.45x
-mvt 7.07 7.06 40,78,43,30 21,64,33,26 1.00x
-Average 27.39 23.58 1.84x
-Geo Mean 4.71 6.95 1.47x
-Table 4. Comparison of GF/s and resources utilization with HARP
-6.4. Latency estimation
-Implementing the post-optimization strategy outlined in Section 5.2 leads to a loss of the lower bound discussed in (Pouget et al., 2024a, b), primarily because we do not incorporate computation-communication overlap modeling into the NLP.
-This omission is deliberate, as including such modeling would significantly increase search time.
-Similarly, we opt against integrating the loop_flatten pragma into our model, despite its potential automatic invocation by the compiler for loops without reductions.
-As detailed in Section 5, we intentionally deactivate it for loops with reductions.
-Although this optimization’s automatic application results in the loss of the lower bound, we have made this choice owing to its rare implementation.
-Our aim is to avoid reliance on an optimization that may not be consistently available.
-In Figure 1, we depict the estimated latency by the NLP of the evaluated designs and juxtapose them with the latencies reported by HLS.
-Our post-optimization efforts reveal that most latencies estimated by NLP exceed those reported by HLS, with the exception of CNN with and without tree reduction.
-In the case of CNN (the last two dots), the pragmas are implemented as anticipated, but the lower bound of the computed unrolled part is not sufficiently tight.
-Figure 1. Comparison of the latency reported by HLS with the estimated latency obtained through NLP
-6.5. Scalability of the NLP
-Among the 42 NLP evaluations (including those with pessimistic resource reuse), only 3 timed out after 4 hours.
-These timeouts were specifically encountered during the bicg and atax kernels with tree reduction, and the gemver kernel without tree reduction.
-Throughout the search process, the NLP solver BARON maintains two bounds: a lower bound, which may be unattainable, and the lowest achievable bound discovered thus far.
-If the solver completes before reaching the timeout, these two bounds are identical.
-In the cases of atax and bicg, these two bounds were within 1% of each other, while for gemver, the difference was 20%.
-The average solving time was 33 minutes, while the geometric mean stood at 2.94 minutes.
-Out of 42 instances, 35 finished within an hour.
+Kernel HARP Ours HARP Sisyphus
+atax 1.72 1.96 78,52,49,50 38,65,53,33 1.14x
+bicg 0.92 1.96 75,25,35,36 38,65,53,34 2.13x
+gemm 125.59 88.61 29,80,55,40 32,48,24,17 0.71x
+gemver 1.66 9.42 29,28,35,19 30,64,65,33 5.69x
+mvt 7.07 8.24 40,78,43,30 53,43,49,30 1.17x
+Average 27.39 22.04 2.16x
+Geo Mean 4.71 7.66 1.62x
+Table 6. Comparison with HARP Using Vitis HLS
+6.4. Single SLR Onboard Evaluation
+GF/s RU: BRAM,DSP,LUT(K),FF(K) Perf. Impr.
+Kernel Ours Ours AutoDSE
+2mm 44.32 312,549,83,94 353.5,963,284,292 126.6x
+3mm 29.89 411,980,108,123 470,1117,281,306 76.6x
+atax 1.03 252,169,118,79 630.5,452,170,212 4.47x
+bicg 1.03 253,171,131,86 867.5,196,168,217 1.88x
+Table 7. Comparison of Onboard Evaluation for 1 SLR
+The resource utilization (RU) is detailed in the thousands for both LUTs and FFs.
+URAM utilization is excluded as no kernels use it.
+And the last column show the performance improvement we achieve with Sisyphus over AutoDSE.
+For the kernels evaluated on board, we can see that our generated design is both faster and consumes fewer resources compared to AutoDSE.
+6.5. Latency Estimation and Scalability
+Implementing the post-optimization strategy from Section 5 results in a loss of the lower bound discussed in (Pouget et al., 2024a, b), mainly because we exclude communication overlap modeling between loop bodies in the NLP to avoid significantly increasing search time.
+We also do not incorporate the loop_flatten pragma, even if the compiler may automatically apply it for loops without reductions, as we deactivate it for reduction loops.
+Although this leads to a lower bound loss, we prefer not to depend on an optimization that may not be consistently implemented.
+The average solving time of the NLP was 455 seconds, while the geometric mean stood at 6.7 seconds.
+Only the kernel 3mm with tree reduction timeout at 4h.
+32/36 NLP’s problem are solve in less than 1 minutes.
+It is important to note that the solver provides a feasible solution early in the process, though it may not be theoretically optimal.
+As the search continues, the solution improves, and once the solver completes, the solution is theoretically optimal.
+Therefore, it is possible to set a timeout to obtain a feasible, but not optimal, solution more quickly.
 7. Related Work
-Design Space Exploration (DSE) methodologies for pragma insertion, such as those mentioned in (Pouget et al., 2024a, b; Sohrabizadeh et al., 2021, 2023, 2022; Ferretti et al., 2018; Zhao et al., 2017; Sohrabizadeh et al., 2023; Zhong et al., 2016; Wu et al., 2023; Liu and Carloni, 2013; Mahapatra and Schafer, 2014; Yu et al., 2018), yield designs with a satisfactory QoR.
-However, they often require extensive computational time, with AutoDSE taking up to a day to complete.
-While users have the option to perform code transformations before conducting DSE, the process typically involves treating each code transformation and pragma insertion as separate optimization problems.
-This decomposition of the problem can result in a loss of QoR because predicting which transformation is necessary to achieve an optimal design is challenging.
-The NLP-DSE method, as explained in (Pouget et al., 2024b, a), employs a non-linear solver to integrate pragmas into the code.
-We build upon the methodology developed by the authors and tailor it to our specific optimization space to determine the pragmas and schedule for the kernel.
-Various code transformations have been devised for CPUs (Bondhugula et al., 2008), GPUs (Verdoolaege et al., 2013), and FPGAs (Pouchet et al., 2013; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018).
+Various code transformations have been devised for CPUs (Bondhugula et al., 2008), GPUs (Verdoolaege et al., 2013), FPGAs (Pouchet et al., 2013b; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) or for all (Zheng et al., 2020).
 While code transformations for CPUs and GPUs yield remarkable results tailored to their respective architectures, they may not be inherently suitable for FPGA targets aimed at achieving high parallelism.
-Pluto (Bondhugula et al., 2008), considered state-of-the-art, performs code transformations including tiling to minimize dependency distance between memory accesses, thereby facilitating data reuse.
-However, in our scenario, reducing dependency distance could potentially constrain parallelization efforts.
 Regarding (Zhao et al., 2022; Zhao and Cheng, 2021), they employ Pluto on different scopes of the kernel for code transformation, yet its pragma insertion capabilities are limited, making it incomparable to our work.
-Conversely, (Pouchet et al., 2013; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) pursue a distinct objective from ours.
-(Pouchet et al., 2013) aims to minimize communication between off-chip and on-chip, which results in better QoR than ours for memory-bound kernels.
+Conversely, (Pouchet et al., 2013b; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) pursue a distinct objective from ours.
+(Pouchet et al., 2013b) aims to minimize communication between off-chip and on-chip, which results in better QoR than ours for memory-bound kernels.
 The works (Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) concentrate on code transformations aimed at enhancing pipelining techniques.
 However, these objectives may not be suitable for computation-bound kernels requiring high levels of parallelization.
-The (Ye et al., 2022; Lai et al., 2019; Huang et al., 2021; Chen et al., 2024; Ye et al., 2024) compilers undertake code transformation and pragma insertion.
-However, these transformations are somewhat limited, mostly encompassing heuristic modifications based on loop properties.
-Moreover, the scope of pragma insertion is more limited compared to our proposed approach.
+Ansor (Zheng et al., 2020) achieves impressive results for CPU and GPU architectures; however, there are notable differences with our approach.
+Ansor requires multiple evaluations to perform its design-space exploration (DSE) and update its cost model, which accelerates the DSE process but is not scalable for FPGA, where a single evaluation can take minutes or even hours.
+Furthermore, Ansor’s approach to code transformations and pragma insertions (which are more tailored for CPU/GPU) addresses these problems separately, potentially resulting in suboptimal designs.
 The choice of tile sizes significantly impacts the QoR.
@@
 On the other hand, (Li et al., 2021) investigates tiles size selection for Convolutional Neural Networks (CNNs) with three-level CPU caching.
+8. Limitations
+Sisyphus is designed to optimize affine loop nests.
+For applications with non-affine code, we can partition the code into affine and non-affine sections.
+Sisyphus handles affine regions effectively, while tools like AutoDSE (Sohrabizadeh et al., 2021) or HARP (Sohrabizadeh et al., 2023) can optimize non-affine parts.
+The optimized sections are then integrated using a dataflow model, either with FIFOs for data exchange or by employing TAPA (Guo et al., 2023), which facilitates dataflow programming.
+This combined approach ensures that both affine and non-affine portions of the application are handled in an optimal manner, achieving a balance between different types of optimizations and seamlessly integrating them for improved overall performance.
 9. Conclusion
-This paper addresses challenges in optimizing HLS via code transformation and pragma insertion.
-While HLS and source-to-source compilers accelerate hardware design, achieving high-quality results often demands manual code intervention, prone to errors and time-consuming.
-To address this, Sisyphus is introduced, automating code transformation and pragma insertion for HLS into a single optimization problem.
-Using Non-Linear Programming (NLP), Sisyphus determines code transformations and pragma placements simultaneously, ensuring optimized performance within resource constraints.
+In this paper, we introduced Sisyphus, a unified framework that streamlines hardware design through high-level synthesis.
+By integrating code transformation, pragma insertion, and tile size selection into a single optimization, Sisyphus addresses challenges in existing HLS workflows, which often rely on manual and disjointed methods.
+Sisyphus defines a design space in the form of a template.
+Using NLP, it efficiently explores this space, quickly identifying transformations and pragma configurations to achieve designs with good QoR.
+Our evaluation demonstrates that Sisyphus outperforms AutoDSE, NLP-DSE, and ScaleHLS in terms of QoR.

### v5

@@
-Enhancing High-Level Synthesis with Automated Pragma Insertion and Code Transformation Framework
+A Unified Framework for Automated Code Transformation and Pragma Insertion
 Abstract.
@@
 It is too difficult to navigate this way, limits the exploration of potential optimizations, and complicates the design generation process.
-To tackle this obstacle we propose Sisyphus, a unified framework that automates code transformation, pragma insertion, and tile size selection within a single optimization problem.
+To tackle this obstacle, we propose Sisyphus, a unified framework that automates code transformation, pragma insertion, and tile size selection within a common optimization framework.
 By leveraging Nonlinear Programming, our approach efficiently explores the vast design space of regular loop-based kernels, automatically selecting loop transformations and pragmas that minimize latency.
@@
 HLS, code transformation, pragma insertion, non-linear problem
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; March xx–xx, 2025; California, USAisbn: 978-1-4503-XXXX-X/18/06
+journalyear: 2025copyright: rightsretainedconference: Proceedings of the 2025 ACM/SIGDA International Symposium on Field Programmable Gate Arrays; February 27–March 1, 2025; Monterey, CA, USAbooktitle: Proceedings of the 2025 ACM/SIGDA International Symposium on Field Programmable Gate Arrays (FPGA ’25), February 27–March 1, 2025, Monterey, CA, USAdoi: 10.1145/3706628.3708873isbn: 979-8-4007-1396-5/25/02ccs: Hardware High-level and register-transfer level synthesisccs: Software and its engineering Compilers
 1. Introduction
 High-Level Synthesis (HLS) compilers and source-to-source compilers are indispensable tools in accelerating hardware design by automating the translation of high-level programming languages like C/C++ or Python into hardware descriptions.
-They offer notable advantages, such as reduced development time and enhanced performance for hardware designs (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020; Zhang et al., 2024).
-However, achieving a good quality of results (QoR) often requires manual code transformations and pragma insertions, which can be facilitated with either Design Space Exploration (DSE) (Pouget et al., 2024a, b; Sohrabizadeh et al., 2021, 2023, 2022; Ferretti et al., 2018; Zhao et al., 2017; Sohrabizadeh et al., 2023) or source-to-source compilers (Lai et al., 2019; Ye et al., 2022, 2024; Zhang et al., 2024) to guide the synthesis process.
+They offer notable advantages, such as reduced development time and enhanced performance for hardware designs (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020; Zhang et al., 2024; Cong et al., 2022).
+However, achieving a good quality of results (QoR) often requires manual code transformations and pragma insertions, which can be facilitated with either Design Space Exploration (DSE) (Pouget et al., 2024, 2025; Sohrabizadeh et al., 2022b, 2023, a; Ferretti et al., 2018; Zhao et al., 2017; Sohrabizadeh et al., 2023) or source-to-source compilers (Lai et al., 2019; Ye et al., 2022, 2024; Zhang et al., 2024) to guide the synthesis process.
 These pragmas aid in optimizing the generated hardware code.
@@
 Additionally, code transformations and pragma insertion are often treated as separate processes.
-Transformations such as loop tiling, splitting, and permutation are typically performed in advance or as part of a pre-processing stage, while pragmas for optimizations like loop unrolling, pipelining, and array partitioning are applied later.
+Transformations such as loop tiling, splitting, and permutation are typically performed in advance or as part of a pre-processing stage, followed by pragma insertion for optimization, like loop unrolling, pipelining and array partitioning.
 This separation introduces inefficiencies, as transformations can affect the effectiveness of pragmas, and vice versa.
@@
 This method facilitates rapid exploration of the entire theoretical space.
-We have developed an analytical model that integrates considerations of latency and resource utilization, building upon previous research (Pouget et al., 2024b, a).
+We have developed an analytical model that integrates considerations of latency and resource utilization, building upon previous research (Pouget et al., 2025, 2024).
 Our focus lies specifically on regular loop-based kernels (PolyBench, [n.
@@
 In summary, we introduce the following contributions:
-  • A novel template that integrates code transformation, pragma insertion, and tiles size selection for data caching by consolidating these tasks into a single optimization problem.
+  • A novel optimization template that integrates code transformation, pragma insertion, and tiles size selection for data caching by consolidating these tasks into a single optimization problem.
 This approach not only simplifies the search space but also ensures that only legal transformations are considered.
   • A novel NLP-based approach specifically designed to explore the joint design space of regular loop-based kernels.
-Unlike traditional methods that involve time-consuming DSE, our approach provides solutions within seconds or minutes.
-Additionally, the NLP is user-friendly, allowing users to easily configure parameters, such as on-chip buffer size, to meet the specific requirements of the kernel.
+Unlike traditional methods that involve time-consuming DSE, our approach provides solutions within minutes.
+Plus, the NLP is user-friendly, allowing users to easily configure parameters, such as on-chip buffer size, to meet the specific requirements of the kernel.
   • Our framework acts as a comprehensive, fully automated system, offering end-to-end functionality.
-With it, we can conduct thorough evaluations and attain QoR designs that are comparable or superior to those achieved by AutoDSE, NLP-DSE, ScaleHLS and HARP, all without the necessity of DSE.
-Furthermore, Sisyphus accurately identifies the appropriate transformation, even in cases where no transformation is required.
+With it, we can conduct thorough evaluations and attain QoR designs that are comparable or superior to those achieved by AutoDSE, NLP-DSE, ScaleHLS and HARP – all without requiring DSE to perform multiple HLS compiler syntheses.
+Furthermore, Sisyphus identifies the correct transformation, even when no transformation is needed.
 The paper is organized into the following sections.
-Section 2 lays out the motivation behind our approach and the proposed solution.
-Following that, Section 3 presents the range of code transformation and pragma insertion we consider.
-In Section 4, we introduce a nonlinear formulation based on this model to automatically discover schedule and pragma configurations through NLP optimization.
+Section 2 explains our approach and solution.
+Following that, Section 3 covers code transformation and pragma insertion.
+Section 4 introduces a nonlinear formulation based on this model to automatically discover schedule and pragma configurations through NLP optimization.
 Section 5 elaborates on our post-optimizations.
@@
 2. Background and Motivation
-2.1. Space to Explore
+2.1. Design Space to Explore
 The exploration space for optimizing an HLS design includes pragma insertion, code transformations, and data caching, all within the limits of resource constraints.
@@
 For example, unrolling pragmas affect array partitioning, which in turn influences BRAM utilization.
-Similarly, the scheduling and pragmas impact the potential for on-chip data caching, while the choice of schedule dictates how loops should be unrolled based on their characteristics, such as whether they are reduction loops.
+Similarly, loop scheduling and pragmas impact the potential for on-chip data caching, while the choice of schedule dictates how loops should be unrolled based on their characteristics, such as whether they are reduction loops.
 A reduction loop is a loop where an operation, like addition or multiplication, accumulates values across iterations to produce a single output, often by reducing a multi-element array to a scalar (e.g., loop L2 in Listing 1).
 The choice of schedule for reduction loops, as opposed to non-reduction loops (e.g., L0, L1, and L3 in Listing 1), influences unrolling strategies and overall optimization because of the dependencies inherent to the loop’s characteristics.
-Consequently, selecting one element impacts the others, often requiring backtracking when the spaces are treated separately.
-This interconnection adds complexity to the exploration process and necessitates ongoing verification to ensure the legality of transformations.
+Selecting one optimization affects others, often requiring backtracking when spaces are handled separately.
+This interdependence complicates exploration and demands constant verification of transformation legality.
 1for (i = 0; i < 200; i++) { // L0
@@
 75store_C(C, C_off_chip);
-Listing 1: Original Code
+Listing 1: Original gemm Code
 1load_C(C, C_off_chip);
@@
 33store_C(C, C_off_chip);
-Listing 2: Result found using the NLP with the constraints:
-DSP = 6,840 On-chip mem = 7.2 MB
-2.2. Limitation of the Current DSE
+Listing 2: Result found for gemm kernel using the NLP with the constraints: DSP = 6,840 On-chip mem = 7.2 MB
+2.2. Limitation of the Current DSE Methods
 Several frameworks (Lai et al., 2019; Ye et al., 2022; Zhao et al., 2022; Ye et al., 2024; Pouchet et al., 2013a; Zhang et al., 2024; Chen et al., 2024) enable code transformations and pragma insertion, but have limited exploration spaces (Ye et al., 2022; Zhao et al., 2022; Ye et al., 2024; Pouchet et al., 2013a; Zhang et al., 2024) or lack automatic DSE (Lai et al., 2019; Chen et al., 2024).
 Transformations are often restricted to loop property-based permutations (Lai et al., 2019; Ye et al., 2022, 2024; Pouchet et al., 2013a; Zhang et al., 2024) or rely on Pluto (Bondhugula et al., 2008), which is optimized for CPUs but overlooks FPGA-specific techniques like pipelining and array partitioning.
-Additionally, Pluto minimizes memory transfer from off-chip to on-chip (Comm), which can limit parallelism on FPGAs, unlike our focus on reducing global latency (Lat).
-Meanwhile, other DSE frameworks (Sohrabizadeh et al., 2021, 2023, 2022; Pouget et al., 2024b) explore pragmas insertion for fixed loop orders, but predicting friendly transformations beforehand is challenging.
+Additionally, Pluto minimizes memory transfer from off-chip to on-chip communication (Comm), which can limit parallelism on FPGAs, unlike our focus on reducing global latency (Lat).
+Meanwhile, other DSE frameworks (Sohrabizadeh et al., 2022b, 2023, a; Pouget et al., 2025) explore pragmas insertion for fixed loop orders, but predicting friendly transformations beforehand is challenging.
 Exploring every possible transformation is also impractical for large design spaces.
@@
 We will now examine examples that highlight the limitations of current methods.
-For this, we use two different HLS DSE methods, AutoDSE (Sohrabizadeh et al., 2021) and NLP-DSE (Pouget et al., 2024b, a).
+For this, we use two different HLS DSE methods, AutoDSE (Sohrabizadeh et al., 2022b) and NLP-DSE (Pouget et al., 2025, 2024).
 A comparison with HARP (Sohrabizadeh et al., 2023) and ScaleHLS (Ye et al., 2022) will be presented in Section 6.
-AutoDSE treats the compiler Merlin (Xilinx, 2024) as a black box, adjusting pragmas based on identified bottlenecks from previous iterations.
+AutoDSE treats the source-to-source compiler Merlin (Xilinx, 2024) and the associated HLS tools as a black box, adjusting pragmas based on the identified bottlenecks from previous iterations.
 In contrast, NLP-DSE employs Nonlinear Programming DSE, utilizing a lower bound-based objective function to achieve high QoR within a short timeframe.
-Both of this DSE use the AMD/Xilinx source-to-source compiler Merlin (Xilinx, 2024).
+Both of this DSE use the AMD source-to-source compiler Merlin (Xilinx, 2024).
 The compiler employs different code transformations such as strip-mining via the TILE pragma or for partial loop unrolling.
 It typically avoids permutations, except when partially unrolling the two innermost loops, in which case the compiler strip-mines these loops and applies permutations to the resulting fully unrolled ones.
-A more detailed description of the Merlin compiler can be found in (Pouget et al., 2024b).
+A more detailed description of this compiler can be found in (Pouget et al., 2025).
 In the following examples, we use the HLS compiler Vitis 2023.2 with the unsafe-math option disabled.
@@
 2.3. Overview of Sisyphus
-To overcome these limitations, we introduce Sisyphus, a solution tailored for affine programs, where loop bounds and conditionals are affine functions of surrounding loop iterators and parameters (commonly referred to as Static Control Parts in the Polyhedral Model) (Girbal et al., 2006; Feautrier, 1988).
-Our objective is to define a search space that allows for the selection of various parameters, such as pragma insertion, loop ordering and tiling, within a single optimization problem.
-This approach avoids the need for backtracking or handling disjoint optimization problems.
-Our framework efficiently and rapidly explores this large design space, overcoming the limitations of other frameworks that either restrict the search space too much or explore a broad space with a DSE that only partially covers it.
-Most importantly, it guarantees that only valid code transformations are considered.
-To accomplish this, we propose a two-step process: first, the creation of a space tailored to each kernel, followed by the exploration of that space.
+We introduce Sisyphus, tailored for affine programs with loop bounds and conditionals as affine functions (Girbal et al., 2006; Feautrier, 1988).
+It defines a unified search space for parameters like pragma insertion, loop ordering, and tiling, avoiding backtracking or disjoint optimizations.
+Sisyphus efficiently explores this design space, overcoming limitations of frameworks that either overly restrict or partially explore it.
+Crucially, it ensures only valid transformations.
+The process involves creating a kernel-specific space and exploring it.
 The first step involves creating a template with fully distributed code that includes three levels of loops, each targeting specific optimizations.
@@
 As a result, each original loop can be utilized across all three levels.
-For instance, in Listing 1, the loop $k$ (L2) is used in level 0 alongside loop L5 and in level 2 with loop L11 in Listing LABEL:lst:small_mem.
+For instance, in Listing 1, the loop $k$ (L2) is used in level 0 (L5) and in level 2 (L11) in Listing LABEL:lst:small_mem.
 Specifically, we have an innermost level for fine-grained unrolling, a level dedicated to pipelining, and another level focused on on-chip data caching based on available resources.
 As demonstrated in Listings LABEL:lst:small_mem and 2, depending on the resource constraints specified by the user, we obtain two different code versions.
-For instance, a reduced on-chip memory requirement necessitates the partial transfer of arrays $A$ and $B$ (line 17, 18 in Listing LABEL:lst:small_mem), leading to the use of a smaller on-chip buffer size that meets the user’s constraints.
+For instance, a reduced on-chip memory requirement necessitates the partial transfer of arrays $A$ and $B$ (lines 17, 18 in Listing LABEL:lst:small_mem), leading to the use of a smaller on-chip buffer size that meets the user’s constraints.
 Once this template is generated, it establishes a search space where the unknowns consist of the loop bounds, loop ordering, buffer sizes for on-chip arrays, and the locations in the code where data transfers take place.
-Although other DSE methods can explore this space, we chose to use a cost model formulated as a Nonlinear Programming (NLP) problem.
+Although other DSE methods can explore this space, we chose to use a cost model formulated as a Nonlinear Programming (NLP) problem as in NLP-DSE (Pouget et al., 2025).
 This approach allows us to efficiently navigate the theoretical space within seconds or minutes, leveraging accurate modeling made possible by compile-time analysis for affine programs.
-It is important to emphasize that exhaustive exploration using a simple cost model or synthesizing each design with HLS, as done in AutoDSE, would be impractical due to the sheer size of the search space.
+It is important to emphasize that exhaustive exploration using a simple cost model or synthesizing each design with HLS (Sohrabizadeh et al., 2022b) would be impractical due to the sheer size of the search space.
 In contrast, the NLP approach enables us to explore this space quickly and efficiently, allowing for a more comprehensive investigation than traditional methods.
+Sisyphus outperforms NLP-DSE (Pouget et al., 2025, 2024) by exploring a larger optimization space, including tiling and loop permutations, unlocking performance unattainable by NLP-DSE, which requires separate runs for each transformation.
+Automating the entire process, Sisyphus reduces exploration time and delivers superior results.
+It generates Vitis HLS pragmas with broader parameter support, like array partitioning, enhancing accuracy and design space coverage.
+Its high model precision enables the generation of at most one or two optimized designs, avoiding NLP-DSE’s extensive iterations.
 For gemm, Sisyphus significantly enhances parallelism by splitting and permuting loops without relying on the reduction loop.
@@
 This small unrolling, selected by the NLP, deepens the pipeline, improving the quality of results (QoR) even though the reduction must remain sequential.
-For the CNN, the NLP selects all parameters within the design space, including loop order after exploring 6!
-possibilities, data transfer points, on-chip array size, and the level of parallelism.
-This comprehensive analysis takes just 13 minutes and yields a throughput of 341 GF/s.
-Thanks to a well-defined space that only considers legal transformations and exploration guided by a cost model formulated as an NLP, our framework is able to select the theoretical optimum in the entire space in a matter of seconds or minutes and generate a design.
+For the CNN, the NLP selects all parameters in the design space, including loop order (exploring 6!
+possibilities), on-chip array size, parallelism levels, etc.
+This analysis completes in 13 minutes, achieving a throughput of 341 GF/s.
+Thanks to a well-defined space that considers only legal transformations and exploration guided by a cost model formulated as an NLP, our framework can identify the theoretical optimum across the entire space and generate a design within seconds or minutes.
 3. Unified Space
@@
 We illustrate our code transformation process step by step using the example from Listings 1 to 2.
-Input  Sisyphus takes an affine C/C++ file as input and automatically extracts a polyhedral representation of the code using PoCC (PoCC, [n.
-d.]).
-This representation allows us to gather all relevant information from the code, such as loop order, loop bounds, array sizes, and more.
-We will use this information to create the template and to automatically generate the NLP discussed in Section 4.
-Our framework will be particularly effective for codes that are fully distributable and fully permutable.
-However, in cases where the transformations described later are not legally possible, our framework still functions but restricts the solution space more.
-Maximal Distribution  To kickstart the process, we begin by thoroughly distributing the program, with the goal of maximizing the distance between each statement.
-This distribution strategy creates ample opportunities for parallelization by maximizing the distance between dependencies (reverse of fusion).
-Typically, this results in one statement per loop body with perfectly nested loops.
-To achieve this, we employ ISCC (Verdoolaege, 2011).
-It allows us to explore various schedule distributions and validate the legality of transformations by ensuring dependency constraints are preserved.
-The code presented in Listing LABEL:lst:distributed displays the fully distributed version of the code from Listing 1.
+Input  Sisyphus takes an affine C/C++ file as input and extracts a polyhedral representation using PoCC (PoCC, [n.
+d.]), gathering information like loop order, bounds, and array sizes.
+This data is used to create the template and generate the NLP.
+The framework is most effective for fully distributable and permutable codes but still works with restricted solution space when transformations are not fully legal.
+Maximal Distribution  To initiate the process, we distribute the program to maximize the distance between statements, promoting parallelization by reducing dependency overlap.
+Typically, this yields one statement per loop body within perfectly nested loops.
+Using ISCC (Verdoolaege, 2011), we explore scheduling options and verify transformation legality by preserving dependencies.
+Listing LABEL:lst:distributed shows the fully distributed version of the code from Listing 1.
 Strip-mining  After maximal distribution, we apply strip-mining to each loop twice.
@@
 Flattening the loops at Level 1 into a single loop and then pipelining the resulting loop introduces significant design complexity and leads to excessively long synthesis times.
-Thus, if a loop at Level 1 is pipelined, it implies that the trip counts of the other loops at the same level are reduced to one, ensuring manageable design complexity and synthesis times.
-For instance, in Listing 2, the loop $j$ (L8) is pipelined for the second statement (S1).
+Thus, if a loop at Level 1 is pipelined, it implies that the trip counts of the other loops at the same level are reduced to one.
+For instance, in Listing 2, the loop $j$ (L8) is pipelined for the statement S1.
 This results in the loops $k$ and $i$ (L10 and L11 in Listing 3 ) having a trip count of 1, enabling them to be eliminated from the code.
@@
 In Listings LABEL:lst:small_mem and 2, we can see that the NLP selected two different permutations for the statement S1 (loops L5, L6, and L7).
+Fusion  We have opted not to incorporate fusion into our model for several reasons.
+Firstly, our objective is to maximize parallelization for computation-bound kernels.
+Fusion may reduce or minimize the dependency distances between statements, limiting potential parallelization opportunities.
+Additionally, since our focus is on computation-bound kernels, the primary bottleneck lies in computation.
+Therefore, we are willing to incur a minor cost, even if it involves transferring a tiled array multiple times.
+Overal Template  In the context of our example, these transformations yield the code presented in Listing 3, with each level corresponding to specific transformations.
+The iterator of the loop retains its original name, with an added number indicating the level.
+For example, loop L14 in Listing 3 refers to loop L2 in Listing 1 at level 2 of the strip mining.
+Notably, our framework is capable of preserving the original loop order, allowing it to retain the existing loop structure if the code is already optimized.
+Design Space  The challenge is to determine loop trip counts (e.g., $J0\_S0$) and level 0 loop order while meeting DSP and on-chip memory constraints.
+Hence, the trip counts corresponding to fully unrolled loops must be constrained to avoid over-utilization of DSP resources.
+The on-chip buffers size must not exceed the on-chip memory capacity.
+Consequently, for a large problem size, the outermost level will enable control over the buffer sizes.
 1int perm_S0[2] // Permutation order (0 for i0, 1 for j0)
@@
 […]
-Listing 3: Template
-Fusion  We have opted not to incorporate fusion into our model for several reasons.
-Firstly, our objective is to maximize parallelization for computation-bound kernels.
-Fusion may reduce or minimize the dependency distances between statements, limiting potential parallelization opportunities.
-Additionally, since our focus is on computation-bound kernels, the primary bottleneck lies in computation.
-Therefore, we are willing to incur a minor cost, even if it involves transferring a tiled array multiple times.
-Template  In the context of our example, these transformations yield the code presented in Listing 3, with each level corresponding to specific transformations.
-The iterator of the loop retains its original name, with an added number indicating the level.
-For example, loop L14 in Listing 3 refers to loop L2 in Listing 1 at level 2 of the strip mining.
-Notably, our framework is capable of preserving the original loop order, allowing it to retain the existing loop structure if the code is already optimized.
-Space  Therefore, the challenge involves determining the trip counts of each loop (e.g., $J0\_S0$) and the loop order of the level 0.
-Additionally, the design must adhere to resource constraints, including the usage of DSP and on-chip memory.
-Hence, the trip counts corresponding to fully unrolled loops must be constrained to avoid over-utilization of DSP resources.
-The on-chip buffers size must not exceed the on-chip memory capacity.
-Consequently, for a large problem size, the outermost level will enable control over the buffer sizes.
+Listing 3: Template automatically generated for the gemm input code (Listing 1)
 4. NLP
-Presented here is a comprehensive set of constraints and variables employed in a nonlinear program aimed at discovering the theoretical solution space outlined in Section 3.
-We employ the methodology proposed by (Pouget et al., 2024b, a), adapted to our specific context.
-Similar to their approach, we establish an estimation of the latency by assuming optimistic DSP utilization, assuming perfect resource reuse.
-However, we afford the user the flexibility to adjust the DSP limit or opt for a pessimistic DSP utilization scenario, where no reuse between statements is considered.
-Moreover, users have the ability to adjust the size of on-chip memory, define the maximum number of array partitions, and customize the latency and resources allocated for each operation.
-This flexibility ensures adaptability across various platforms and compilers.
+We automatically generate constraints and variables for a nonlinear program to explore the theoretical solution space outlined in Section 3, adapting the approach from (Pouget et al., 2025, 2024) to our context.
+Latency is estimated with optimistic DSP utilization, assuming perfect resource reuse, but users can adjust the DSP limit or choose a pessimistic scenario with no reuse.
+Users can also customize on-chip memory size, maximal array partitioning, and latency/resources per operation, ensuring adaptability to different platforms and compilers.
 To gather all the required information for the NLP, we utilize PoCC (PoCC, [n.
@@
 4.1. Variables
-Table 2 delineate the sets, variables, and constants utilized in our NLP formulation.
+Table 2 defines the sets, variables, and constants utilized in our NLP formulation.
 4.2. Constraints
-Now, we explore the precise meaning and implications of each constraint.
-To make this more understandable, we provide examples using the listings presented in the paper.
+We describe each constraint’s meaning and implications with examples from the paper’s listings.
 Trip Count  Equation 1 constrains the trip count of each loop, ensuring that the product of the trip counts equals the original trip count.
@@
 (1) $\forall l\in\mathcal{L},\prod_{v\in\mathcal{V}}TC_{l,v}=TC_{l}$
-Pipeline  Constraints 2, 3, 4, and 5 facilitate the selection of loop pipelining.
-Constraint 2 ensures that each loop has a boolean flag indicating whether the original loop is pipelined.
-For example, in Listings LABEL:lst:small_mem and 2, the statement $S1$ shows that the loop $j1$ (L8) is pipelined, while the original loop is $j$ (L3 in Listing 1), indicating that $loop_{j}\_pip=1$.
-As explained in Section 3, only one loop can be pipelined at each middle level.
-This means the sum of the boolean flags must be less than or equal to 1 (Eq.
-4), and the trip count of all other non-pipelined loops at the middle level must be 1 (Eq.
-3).
-For example, in the template (Listing 3), only loop L9 is pipelined at level 1 of S1 (loops L9, L10, and L11), so $loop_{j}\_pip=1$.
-Consequently, $loop_{i}\_pip=0$ and $loop_{k}\_pip=0$ (Eq.
-4), and Eq.
-3 indicates that the trip counts of loops L10 and L11 are 1, allowing their removal.
 Set Description
@@
 $\mathcal{C}_{a_{d}}$ the set of loops iterating the array $a$ at dimension $d$
-$\mathcal{V}$ Level of the strip mining, 0 for coarse-grained/sequential, 1 for pipeline and 2 for fine-grained unrolled
+$\mathcal{V}$ Strip mining level: 0 for coarse-grained/sequential, 1 for pipeline, 2 for fine-grained unrolled
 $AP_{a,d}$ Array Partition for the array $a$ in dimension $d$
@@
 $tc_{l,level}$ TC of the loop $l$ for the level of strip-mining
-$loop_{l}\_UF$ Coarse-grained unroll factor of the loop $l$ at level 0 of the strip-mining
-$loop_{l}\_pip$ Boolean to know if the loop $l$ is pipelined at level 1 of the strip-mining
-$cache_{l}\_arr_{a}$ Boolean to know if the array $a$ is transferred on-chip after the loop $l$ at level 0 of the strip-mining
+$loop_{l}\_UF$ Coarse-grained unroll factor of the loop $l$ at level 0
+$loop_{l}\_pip$ Boolean to know if the loop $l$ is pipelined at level 1
+$cache_{l}\_arr_{a}$ Boolean to know if the array $a$ is transferred on-chip after the loop $l$ at level 0
 $perm_{s}$ ID of the permutation of the statement $s$
 $burst_{a}$ The maximum bitwidth at which the array $a$ can be transferred from off-chip to on-chip
-Table 2. Overview of the set, constant and variables employed in formulating the NLP
+Table 2. Overview of the sets, constants and variables employed in formulating the NLP
+Pipeline  Constraints 25 facilitate the selection of loop pipelining.
+Constraint 2 ensures each loop has a flag indicating if it is pipelined.
+For example, in Listings LABEL:lst:small_mem and 2, $S1$ shows loop $j1$ (L8) is pipelined, while its original loop is $j$ (L3 in Listing 1), making $loop_{j}\_pip=1$.
+As per Section 3, only one loop at level 1 can be pipelined, requiring the sum of pipeline flags must be less than or equal to 1 (Eq.
+4).
+Non-pipelined loops must have a trip count of 1 (Eq.
+3).
+For instance, in Listing 3, only loop L9 is pipelined at level 1 of $S1$, so $loop_{j}\_pip=1$, while $loop_{i}\_pip=0$ and $loop_{k}\_pip=0$.
+Loops L10 and L11 have trip counts of 1 and can be removed.
 Eq.
-5 allows the computation of the initiation interval (II), which is determined based on the loop’s properties and the iteration latency of the reduction operation.
-Specifically, when the loop being pipelined is a reduction loop, this follows the approach outlined in (Pouget et al., 2024b, a).
-In Listings LABEL:lst:small_mem and 2, the pipelined loops are non-reduction loops with no dependencies when iterating, hence $II=1$.
-However, if loop L10 in Listing 3 were pipelined, it would involve a reduction, meaning that at each iteration, C[i][j] needs to be read and written.
-Thus, to read for iteration $k+1$, we must wait for the write to finish at iteration $k$, which takes $IL_{+}$ cycles (latency of the reduction operation).
+5 calculates the initiation interval (II) based on loop properties and reduction operation latency.
+In Listings LABEL:lst:small_mem and 2, pipelined loops are non-reduction loops with no dependencies, giving $II=1$.
+However, if loop L10 in Listing 3 were pipelined, it would involve a reduction requiring $II\geq IL_{+}$, as the write of iteration $k$ must complete before the read of $k+1$.
 (2) $\forall l\in\mathcal{L},loop_{l}\_pip\in\{0,1\}$
@@
 (5) $\forall s\in\mathcal{S},II_{s}=\sum_{l\in\mathcal{L}_{s}}loop_{l}\_pip\times II_{l}$
-Coarse-grained unrolling  We enforce coarse-grained unrolling solely for non-reduction loops (Eq.
-8).
-For example, the loop L3 in Listing LABEL:lst:distributed cannot be coarse-grained unrolled as it is a reduction loop.
-Additionally, the unroll factor (UF) must divide the trip count of the current loop and be less than or equal to it (Eq.
-9 and 6).
-However, as elaborated in Section 3, coarse-grained unrolling is disregarded if there is only one statement within the loop (Eq.
-7).
-This implies that for Listing 1 coarse-grained unrolling is disabled.
+Coarse-grained unrolling  Coarse-grained unrolling applies only to non-reduction loops (Eq.
+6).
+For instance, loop L3 in Listing LABEL:lst:distributed cannot be unrolled as it is a reduction loop.
+The unroll factor (UF) must divide the loop trip count and not exceed it (Eq.
+7 and 8).
+However, as noted in Section 3, coarse-grained unrolling is skipped if the loop has only one statement (Eq.
+9), as in Listing 1.
+(6) $\forall s\in\mathcal{S}_{1},\forall l\in\mathcal{L}_{s}^{red},loop_{l}\_UF=1$
+(7) $\forall l\in\mathcal{L},loop_{l}\_UF\%TC_{l,0}==0$
 (8) $\forall l\in\mathcal{L},loop_{l}\_UF\leq TC_{l,0}$
 (9) $\forall s\in\mathcal{S}_{1},\forall l\in\mathcal{L}_{s},loop_{l}\_UF=1$
-(8) $\forall s\in\mathcal{S}_{1},\forall l\in\mathcal{L}_{s}^{red},loop_{l}\_UF=1$
-(9) $\forall l\in\mathcal{L},loop_{l}\_UF\%TC_{l,0}==0$
-On-chip memory  Constraints 10 and 11 ensure that each array, represented by a boolean (Eq.
-10), is cached on-chip at a single location within the same loop body, as only one boolean can be true for each loop body (Eq.
+On-chip memory  Constraints 10 and 11 ensure each array, represented by a boolean (Eq.
+10), is cached at a single location per loop body, with only one boolean true per loop (Eq.
 11).
-In Listing 3, for instance, array $A$ can be cached either before any loops (line 5) or after L6, L7, or L8 (lines 23, 25, and 27).
-In Listing 2, $A$ is cached on-chip with the on-chip footprint matching the size of the original array (line 2).
-Meanwhile, in Listing LABEL:lst:small_mem, $A$ is cached on-chip at line 18, but with a smaller on-chip footprint compared to the original.
-Based on where each array is transferred, we can compute the on-chip footprint.
-For example, in Listing LABEL:lst:small_mem, array $A$ is transferred below loop L5, which has a trip count of 60, so the footprint of $A$ on-chip is $200\times\frac{240}{60}$.
+In Listing 3, array $A$ can be cached before loops (line 5) or after L6, L7, or L8 (lines 23, 25, 27).
+In Listing 2, $A$ is cached on-chip, matching the original array’s size (line 2).
+In Listing LABEL:lst:small_mem, $A$ is cached on-chip at line 18 with a smaller footprint.
+The on-chip footprint depends on the transfer location.
+For Listing LABEL:lst:small_mem, $A$ is transferred below loop L5 (trip count 48), so its footprint is $200\times\frac{240}{48}$.
 (10) $\forall l\in\mathcal{L},\forall a\in\mathcal{A},cache_{l}\_arr_{a}\in\{0,1\}$
@@
 (12) $\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}cache_{l}\_arr_{a}\times ft\_array_{a}\_loop_{l}\leq Mem$
-Array partitioning  We consider unrolling only if all the data needed for unrolling can be accessed in parallel, i.e., these data are on different on-chip banks thanks to the pragma array_partitioning.
-For each array $a$ and each dimension $d$ of this array, we want the factor of the cyclic array partitioning for this dimension to be greater than the unroll factor (Eq.
-15) to ensure that all data are on different banks.
-If the array is reused, we need to ensure that the array partitioning is optimal for each statement.
-Therefore, we force the array partitioning to be a multiple of each unroll factor that iterates through this dimension (Eq.
+Array partitioning  Unrolling is applied only if all required data can be accessed in parallel, ensured by pragma array_partitioning.
+For each array $a$ and dimension $d$, the cyclic partitioning factor must exceed the unroll factor (Eq.
+13) to ensure data is on separate banks.
+If the array is reused, partitioning must align with each unroll factor iterating through the dimension (Eq.
 14).
-$AP_{a,d}$ gives the cyclic array partitioning factor for array $a$ in dimension $d$ that we use in the generated code.
-Eq.
-13 assures that the array partitioning per array (product of the partitioning factors per dimension) is less than or equal to the limit fixed by the user.
-For example, in Listing 2, $C$ is unrolled 200 times in the first dimension (lines 11, 12) and 4 times in the second dimension (lines 13, 14) for the first loop body, and 200 times in the first dimension (lines 25, 26) for the second loop body.
-Thus, we have $AP_{C,1}=200$ and $AP_{C,2}=4$, and $200\times 4=800\leq max_{part}$, where $max_{part}$ is 1,024 in our case.
+$AP_{a,d}$ represents the cyclic partitioning factor used in the code, with the total partitioning per array constrained by the user-defined limit (Eq.
+15).
+For example, in Listing 2, $C$ is unrolled 200 times in the first dimension (lines 11, 12) and 4 times in the second (lines 13, 14) for the first loop body, and 200 times in the first dimension (lines 25, 26) for the second.
+This gives $AP_{C,1}=200$, $AP_{C,2}=4$, and $200\times 4=800\leq max_{part}$, with $max_{part}=1,024$.
+(13) $\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}\geq TC_{l,2}$
+(14) $\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}\%TC_{l,2}==0$
 (15) $\forall a\in\mathcal{A},\prod_{d\in\mathbb{N}}AP_{a,d}\leq max_{part}$
-(14) $\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}\%TC_{l,2}==0$
-(15) $\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}\geq TC_{l,2}$
-DSP utilization  Finally, we need to manage DSP utilization.
-Given the challenges in predicting DSP reuse, we use a boolean flag, $reuse_{opt}$, set by the user to determine the level of DSP reuse.
-If $reuse_{opt}=1$, we assume an optimistic reuse scenario: DSPs used in one loop body can be directly reused by subsequent loop bodies if the operations are identical.
-However, if the operations differ, as discussed in (Pouget et al., 2024b, a), no reuse is assumed.
-Therefore, the total DSP usage is calculated as the sum of the maximum DSPs required for each operation across all loop bodies (Eq.
+DSP utilization  To manage DSP utilization, we use a user-defined flag, $reuse_{opt}$.
+If $reuse_{opt}=1$, optimistic reuse is assumed: identical operations reuse DSPs across loop bodies, with total usage given by the maximum DSPs per operation (Eq.
 16).
-If $reuse_{opt}=0$, we assume no reuse between different loop bodies, and we sum the DSP requirements for each statement (Eq.
+However, if the operations differ, as discussed in (Pouget et al., 2025, 2024), no reuse is assumed.
+If $reuse_{opt}=0$, no reuse is assumed, and DSPs are summed for each statement (Eq.
 17).
-Based on the value of $reuse_{opt}$, the total DSP usage must be less than or equal to the available DSPs (Eq.
+In both cases, total usage must not exceed available DSPs (Eq.
 18 and 19).
-For instance, if a multiplication operation (*) uses 3 DSPs and an addition operation ($+$) uses 2 DSPs, in Listing 2, S0 requires $200\times 4\times 3$ DSPs for multiplications, and S1 requires $200\times 4\times 3+200\times 4\times 3+200\times 4\times 2$ DSPs for two multiplications and one addition.
-With optimistic reuse for the same operation (e.g., multiplication), the DSP usage is calculated as: $\max(200\times 4\times 3,2\times 200\times 4\times 3)+200\times 4\times 2$.
-With pessimistic reuse, the DSP usage is: $200\times 4\times 3+2\times(200\times 4\times 3)+200\times 4\times 2$.
+If a multiplication (*) uses 3 DSPs and an addition ($+$) uses 2 DSPs, in Listing 2, S0 needs $200\times 4\times 3$ DSPs for multiplications, while S1 requires $200\times 4\times 3+200\times 4\times 3+200\times 4\times 2$ DSPs for two multiplications and one addition.
+With optimistic reuse, DSP usage for multiplication is $\max(200\times 4\times 3,2\times 200\times 4\times 3)+200\times 4\times 2$, while pessimistic reuse results in $200\times 4\times 3+2\times(200\times 4\times 3)+200\times 4\times 2$.
 (16) $DSPs\_used_{opt}=\sum_{op\in\{+,-,*,/\}}\max_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})$
@@
 4.3. Objective Function
-We structure the objective function akin to the approach described in (Pouget et al., 2024b, a), tailoring it to meet the specific needs of our problem.
+We use the objective function similar to the approach described in (Pouget et al., 2025, 2024), tailoring it to meet the specific needs of our problem.
 Here, $Lat^{s}_{2}$ corresponds to the fine-grained unrolled level (level 2) for the statement $s$, $Lat^{s}_{1}$ denotes the pipeline tile that incorporates the fine-grained unrolled level (level 1), and recursively, $Lat_{0}$ encompasses $Lat_{1}$, enabling coarse-grained unrolling and facilitating on-chip data caching (level 0).
@@
 Therefore, the latency at a given level is calculated as the array footprint divided by the burst size (bitwidth used for transferring the array), multiplied by the trip count of the loop that iterates over the function responsible for the transfer, based on the selected permutation of level 0.
-Additionally, if an array is fully transferred on-chip for a statement $s$ and reused by a subsequent statement $s^{’}$, no further transfers are needed.
-The burst size for an array must be a divisor of its last dimension.
-For instance, if we have an array $float\text{ }A[512][20]$, the maximum burst size of this array $A$ is 128 bits, since the last dimension, $20\times 32$ bits, can be divided by the largest power of 2, which is 128.
+Additionally, if an array is fully transferred on-chip for a statement $s$ and reused by a subsequent statement $s{’}$ ’, no further transfers are needed.
+The burst size for an array must divide its last dimension.
+For instance, if we have an array $float\text{ }A[512][20]$, the maximum burst size for this array is $128$ bits.
+This is because $128$ is the largest power of $2$ that can evenly divide the size of the last dimension ($20\times 32$ bits).
 $\left.\begin{cases}&Lat^{s}_{2}=IL_{par}+IL_{seq}\times\prod_{l\in\mathcal{L}^{red}}\times\log_{2}(TC_{l,2})\\ &Lat^{s}_{1}=Lat^{s}_{2}+II\times(TC_{l,1}-1)\\ &Lat^{s}_{0}=\prod_{l\in\mathcal{L}}\frac{TC_{l,0}}{loop_{l}\_UF}\times Lat^{s}_{1}\\ &L^{s}_{mem}=\sum_{l\in\mathcal{L}}\max_{a\in\mathcal{A}}(cache_{l}\_arr_{a}\times ft\_array_{a}\_loop_{l}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\times TCs(perm_{s})/burst_{a})\\ &Lat_{s}=Lat^{s}_{0}+L^{s}_{mem}\end{cases}\right.$
-Finally, the objective function (Eq. 20) is defined as the sum of the latencies of each loop body.
-(20) $obj\_func=\sum_{s\in\mathcal{S}}Lat_{s}$
-5. Post-Optimization and Usage
-The NLP determines the trip count for each loop, array partitioning, and on-chip buffer sizes.
-Functions for loading/storing data between off-chip and on-chip memory are automatically generated with the maximum burst size.
+Finally, the objective function $obj\_func=\sum_{s\in\mathcal{S}}Lat_{s}$ is defined as the sum of the latencies of each loop body.
+5. Post-Optimization and Customization
+The NLP determines the trip counts, array partitioning, and on-chip buffer sizes.
+Data transfer functions between off-chip and on-chip memory are automatically generated with the maximum burst size.
 Further optimizations are also applied to improve QoR.
@@
 For example, instead of using $C[i][i]+=B[i][i]$, we replace it with $C[i][j]=f(C[i][j],B[i][i])$, where the function $f$ ensures compliance with the constraints of the non-constant trip count loop.
-This function returns the computed value if the constraints are satisfied; otherwise, it simply returns the output value.
-Therefore, we employ these techniques to reduce compilation time and achieve designs with a good QoR.
+This function returns the computed value if the constraints are satisfied; otherwise, it simply returns the original output value.
+This technique helps to reduce compilation time and achieve designs with a good QoR.
 Grouping Data Transfers  In the NLP, we already account for parallel transfers of arrays when they are transferred at the same level in the code (e.g., within the same loop).
 However, we do not consider parallel transfers between arrays belonging to different loop bodies.
-In the 2mm example, we first handle the transfer of arrays required for the initial matrix multiplication, only transferring the arrays for the second multiplication after the first is complete.
-However, FPGA DRAM banks enable parallel transfers of different arrays.
-Therefore, we can optimize load and store operations during post-processing to enhance memory transfer overlap.
-Code Transformation for HLS  To simplify the understanding of the HLS compiler, particularly to ensure that the compiler can pipeline loops with the correct initiation interval (II), we simplify the reductions (regardless of whether we use tree reductions) in the fine-grained unrolled part (level 2).
-Therefore, under the pipeline we create a variable to accumulate the reduction, which we then add to the output.
+In our example, we first manage the transfer of arrays required for initializing the matrix, transferring the arrays for multiplication only after the initialization is complete.
+However, FPGA DRAM banks allow parallel transfers of different arrays.
+Consequently, we can optimize load and store operations during post-processing to improve memory transfer overlap.
+Code Transformation for HLS  To help the HLS compiler pipeline loops with the correct initiation interval (II), we simplify reductions at level 2.
+A variable accumulates the reduction within the pipeline, which is then added to the output.
 Additionally, we use the loop_flatten off pragma on all reduction loops above the pipeline.
 Although this adjustment could theoretically improve the QoR, in practice, flattening these loops can complicate the code and potentially increase the initiation interval (II) of the pipeline.
-Usage  Our framework is designed to be explainable and user-friendly, requiring the user to provide only the affine C/C++ code of the kernel to be optimized and the available resources.
+Customization  Our framework is designed to be explainable and user-friendly, requiring the user to provide only the affine C/C++ code of the kernel to be optimized and the available resources.
 To fully benefit developers, our NLP formulation allows the template to be completely customized to the user’s needs.
@@
 For example, in the gemm template (Listing 3), the user can set the permutations for level 0 while allowing the NLP to explore other parameters.
-Additionally, they can specify the on-chip size of matrix A, enabling exploration within the defined constraints.
+Additionally, they can specify the on-chip size of matrix $A$ by directly defining a variable in the NLP, allowing for exploration within the defined constraints.
 6. Evaluation
 The goal of this evaluation is to demonstrate how effectively our approach applies code transformations and inserts hardware directives while respecting user-defined resource constraints.
-We compared our method against four other frameworks: AutoDSE (Sohrabizadeh et al., 2021), NLP-DSE (Pouget et al., 2024b), HARP (Sohrabizadeh et al., 2023), and ScaleHLS (Ye et al., 2022).
+We compared our method against four other frameworks: AutoDSE (Sohrabizadeh et al., 2022b), NLP-DSE (Pouget et al., 2025), HARP (Sohrabizadeh et al., 2023), and ScaleHLS (Ye et al., 2022).
 Since generating a bitstream without significant manual intervention is challenging for some frameworks, we first demonstrate our ability to achieve a high-quality design by utilizing all available resources—ensuring fair comparison with the other frameworks—without considering place and route.
@@
 We use the medium-sized Polybench kernels found in their training set with double-precision floating-point data types, aiming to deploy HARP under optimal conditions.
-As detailed in (Pouget et al., 2024b), HARP’s effectiveness is not universally applicable without the necessary fine-tuning or training, particularly when confronted with diverse problem sizes.
+As detailed in (Pouget et al., 2025), HARP’s effectiveness is not universally applicable without the necessary fine-tuning or training, particularly when confronted with diverse problem sizes.
 We selected these problem sizes to highlight how our method performs in transforming code and inserting pragmas in cases where full unrolling is not feasible.
@@
 When enabling tree reduction, we selected the “funsafe math optimizations” option to allow commutative and associative reduction operators for logarithmic time reductions.
-For our hardware setup, we targeted the Xilinx Alveo U200 device running at 250 MHz.
-We use the commercial solver Gurobi 11.0.0 (Gurobi Optimization, LLC, 2024) with AMPL.
-The experiments were performed on an AMD EPYC 7V13 64-core processor, with 32 GB of memory and 4 cores allocated for each run.
+For our hardware setup, we targeted the Xilinx Alveo U200 device operating at 250 MHz.
+We utilized the commercial solver Gurobi 11.0.0 (Gurobi Optimization, LLC, 2024) with AMPL.
 6.2. Experimental Evaluation
@@
 AutoDSE was run with a 1,000-minute exploration timeout and a 180-minute HLS synthesis timeout.
-For NLP-DSE, we followed the original paper’s parameters.
-HARP was run with the authors’ settings using Vitis 2021.1 and tree reduction enabled.
+For NLP-DSE, we followed the original paper’s parameters (Pouget et al., 2025).
+HARP was run with the authors’ settings (Sohrabizadeh et al., 2023) using Vitis 2021.1 and tree reduction enabled.
 ScaleHLS assumes that all kernel data are already located in on-chip memory.
@@
 However, due to the extensive manual modifications required, we did not extend this comparison to all kernels.
-For Sisyphus evaluation, we set the Gurobi solver timeout to 14,400 seconds (4 hours) per design.
+For Sisyphus evaluation, we set the Gurobi solver timeout to 14,400 seconds per design.
 By default, we assume optimistic resource reuse between non-parallel statements.
@@
 displays the performance improvement of our framework compared to the three different frameworks.
-Across all evaluated kernels, we consistently achieve performance that is comparable to or better than alternatives, except for the atax, and gesummv kernels, where we observe a slowdown of less than 5%.
-For 3mm, symm and trmm we had to apply the constraint indicating no reuse (Eq.
-17), as the constraint with optimistic reuse (Eq.
-16) resulted in kernels with resource over-utilization.
-For Gesummv, ScaleHLS is 1.41x faster.
-However, 91% of our design’s execution time is spent on data transfers.
-ScaleHLS benefits from 512-bit data transfers, which we manually applied for testing.
-Our framework uses 64-bit chunks, the largest bitwidth that divides the array.
-With 512-bit transfers, our framework could achieve 14.50 GF/s.
-For other memory-bound kernels like Atax and Gesummv, the slight performance differences observed with other frameworks are largely due to how Merlin handles memory transfers.
 GF/s Perf. Impr. of ours framework vs.
@@
 Geo Mean 40.13 9.17x 3.47x 16.15x
-Table 3. Throughput (GF/s) Comparison of Sisyphus, AutoDSE, NLP-DSE, and ScaleHLS with Tree Reduction Using Vitis HLS
+Table 3. Throughput (GF/s) Comparison with Tree Reduction Using Vitis HLS
+Across all evaluated kernels, we consistently achieve performance that is comparable to or better than alternatives, except for the atax, and gesummv kernels, where we observe a slowdown of less than 5%.
+For 3mm, symm and trmm we had to apply the constraint indicating no reuse (Eq.
+17), as the constraint with optimistic reuse (Eq.
+16) resulted in kernels with resource over-utilization.
+For Gesummv, ScaleHLS is 1.41x faster.
+However, 91% of our design’s execution time is spent on data transfers.
+ScaleHLS benefits from 512-bit data transfers, which we manually applied for testing.
+Our framework uses 64-bit chunks, the largest bitwidth that divides the array.
+With 512-bit transfers, our framework could achieve 14.50 GF/s.
+For other memory-bound kernels like Atax and Gesummv, the slight performance differences observed with other frameworks are largely due to how Merlin handles memory transfers.
 Sisyphus averages a speedup of 683.91x, 641.46x and 434.78x over AutoDSE, NLP-DSE and Scale-HLS respectively, across the evaluated kernels.
@@
 Table 4 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE) and Sisyphus (ours) without tree reduction.
-Sisyphus achieves an average speedup of 676.03x over AutoDSE and 567.53x over NLP-DSE across the evaluated kernels without tree reduction.
-The geometric mean speedup is 18.50x and 4.80x, respectively.
-On average, the performance improvement without tree reduction is greater than with tree reduction, particularly for gemm (Listing 2) and gemver.
-This is because, without tree reduction, optimizing the reduction loop becomes challenging, making code transformations even more necessary.
-However the difference observed between syrk and syr2k with tree reduction is due to the HLS compiler achieving an II=2 instead of the expected II=1.
-However, without tree reduction, the compiler successfully achieves II=1.
 GF/s Perf. Impr. Ours vs.
@@
 doitgen D,P 54.05 5.46x 2.71x
-gemm D,P 210.61 9.25x 9.57x
+gemm D,P 210.61 9.25x 4.01x
 gemver P 15.01 37.28x 8.73x
@@
 trmm D,P 130.90 7,362.68x 7,333.85x
-Average 133.84 676.03x 567.53x
-Geo Mean 41.62 18.50x 4.80x
-Table 4. Throughput (GF/s) Comparison of Sisyphus, AutoDSE and NLP-DSE without Tree Reduction Using Vitis HLS
+Average 133.84 676.03x 567.10x
+Geo Mean 41.62 18.50x 4.49x
+Table 4. Throughput (GF/s) Comparison without Tree Reduction Using Vitis HLS
+Sisyphus achieves an average speedup of 676.03x over AutoDSE and 567.10x over NLP-DSE across kernels without tree reduction, with geometric means of 18.50x and 4.49x, respectively.
+Performance improvement is higher without tree reduction, especially for gemm (Listing 2) and gemver, as optimizing reduction loops becomes harder, increasing the need for code transformations.
+The difference between syrk and syr2k with tree reduction occurs because the HLS compiler achieves II=2 instead of the expected II=1, while without tree reduction, II=1 is achieved.
 6.3.3. AutoDSE and NLP-DSE for Bert and CNN
-Table 5 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE) and Sisyphus (ours).
-The TR column denotes the status of the tree reduction option.
+Table 5 compares AutoDSE (A-DSE), NLP-DSE (N-DSE), and Sisyphus (ours).
+The TR column shows the tree reduction status.
 For bert_100_64, we observed a minor decrease in QoR, attributed to the NLP selecting a configuration where the pipeline was not applied as expected.
-Code transformation with tiling  The selection of loop order and tile size is critical for determining which array sizes, particularly for large problem size like bert_3072_100 and CNN layers, should be cached on-chip and where they should be transferred.
-In these instances, the NLP aims to enhance data reuse and minimize unnecessary memory transfers while identifying the theoretical optimal balance with parallelism.
 GF/s Perf. Impr. Ours vs.
@@
 Geo Mean 185.32 4.38x 4.69x
-Table 5. Throughput (GF/s) Comparison of Sisyphus, AutoDSE and NLP-DSE Using Vitis HLS
+Table 5. Throughput (GF/s) Comparison Using Vitis HLS
+Code transformation with tiling  The selection of loop order and tile size is critical for determining which array sizes, particularly for large problem size like bert_3072_100 and CNN layers, should be cached on-chip and where they should be transferred.
+In these instances, the NLP aims to enhance data reuse and minimize unnecessary memory transfers while identifying the theoretical optimal balance with parallelism.
 6.3.4. HARP
 In Table 6 we compare the throughput and resource utilization (BRAM, DSP, LUT, FF) achieved with HARP (Sohrabizadeh et al., 2023).
-The enhanced performance seen in both bicg and gemver can be attributed to the same underlying factor observed in the comparison with AutoDSE and NLP-DSE.
-Regarding mvt, HARP produced comparable performance to our own due to their use of the Merlin source-to-source compiler, which facilitates loop permutations when partially unrolling consecutive loops.
-This allows them to find the same schedule that our framework identified.
-However, for Gemm, the design discovered by HARP falls outside our specified parameter space, particularly because of our restriction on maximal array partitioning.
-Notably, it is the only design in the top 10 that outperforms our framework.
 GF/s Resource Utilization (%) Perf. Impr.
@@
 atax 1.72 1.96 78,52,49,50 38,65,53,33 1.14x
-bicg 0.92 1.96 75,25,35,36 38,65,53,34 2.13x
-gemm 125.59 88.61 29,80,55,40 32,48,24,17 0.71x
-gemver 1.66 9.42 29,28,35,19 30,64,65,33 5.69x
-mvt 7.07 8.24 40,78,43,30 53,43,49,30 1.17x
-Average 27.39 22.04 2.16x
-Geo Mean 4.71 7.66 1.62x
+bicg 0.92 1.96 75,25,35,36 38,65,48,31 2.13x
+gemm 125.59 87.86 29,80,55,40 32,48,36,23 0.70x
+gemver 1.66 9.42 29,28,35,19 34,8,31,11 5.69x
+mvt 7.07 8.29 40,78,43,30 53,44,57,30 1.17x
+Average 27.39 20.71 1.45x
+Geo Mean 4.71 6.28 1.33x
 Table 6. Comparison with HARP Using Vitis HLS
+The enhanced performance seen in both bicg and gemver can be attributed to the same underlying factor observed in the comparison with AutoDSE and NLP-DSE.
+Regarding mvt, HARP produced comparable performance to our own due to their use of the Merlin compiler, which facilitates loop permutations when partially unrolling consecutive loops.
+This allows them to find the same schedule that our framework identified.
+However, for gemm, the design discovered by HARP falls outside our specified parameter space, due to our restriction on maximum array partitioning.
 6.4. Single SLR Onboard Evaluation
@@
 Kernel Ours Ours AutoDSE
-2mm 44.32 312,549,83,94 353.5,963,284,292 126.6x
-3mm 29.89 411,980,108,123 470,1117,281,306 76.6x
-atax 1.03 252,169,118,79 630.5,452,170,212 4.47x
-bicg 1.03 253,171,131,86 867.5,196,168,217 1.88x
+2mm 30.57 510,556,213,276 353.5,963,287,292 76.98x
+3mm 29.89 611,984,230,300 470,1117,278,306 72.34x
+atax 1.03 450,173,240,250 630.5,452,170,212 4.63x
+bicg 1.02 451,173,238,265 867.5,196,168,217 1.80x
 Table 7. Comparison of Onboard Evaluation for 1 SLR
@@
 6.5. Latency Estimation and Scalability
-Implementing the post-optimization strategy from Section 5 results in a loss of the lower bound discussed in (Pouget et al., 2024a, b), mainly because we exclude communication overlap modeling between loop bodies in the NLP to avoid significantly increasing search time.
+Implementing the post-optimization strategy from Section 5 results in a loss of the lower bound discussed in (Pouget et al., 2024, 2025), mainly because we exclude communication overlap modeling between loop bodies in the NLP to avoid significantly increasing search time.
 We also do not incorporate the loop_flatten pragma, even if the compiler may automatically apply it for loops without reductions, as we deactivate it for reduction loops.
 Although this leads to a lower bound loss, we prefer not to depend on an optimization that may not be consistently implemented.
+As noted in Section 5, our post-optimization focuses primarily on memory transfer overlap, so our final latency values do not directly reflect the model’s initial predictions.
+However, our model achieves (when predicting the latency of the final design generated) an average prediction error of 8% across evaluated designs, with a 3% error in geometric mean.
+For certain designs, like syrk and syr2k, where the compiler achieves an II=2 rather than the expected II=1, prediction errors are higher (49% and 32%, respectively) due to an optimistic estimation of II.
 The average solving time of the NLP was 455 seconds, while the geometric mean stood at 6.7 seconds.
@@
 For applications with non-affine code, we can partition the code into affine and non-affine sections.
-Sisyphus handles affine regions effectively, while tools like AutoDSE (Sohrabizadeh et al., 2021) or HARP (Sohrabizadeh et al., 2023) can optimize non-affine parts.
+Sisyphus handles affine regions effectively, while tools like AutoDSE (Sohrabizadeh et al., 2022b) or HARP (Sohrabizadeh et al., 2023) can optimize non-affine parts.
 The optimized sections are then integrated using a dataflow model, either with FIFOs for data exchange or by employing TAPA (Guo et al., 2023), which facilitates dataflow programming.
@@
 9. Conclusion
-In this paper, we introduced Sisyphus, a unified framework that streamlines hardware design through high-level synthesis.
+In this paper, we introduced Sisyphus, a unified framework that streamlines hardware design through HLS.
 By integrating code transformation, pragma insertion, and tile size selection into a single optimization, Sisyphus addresses challenges in existing HLS workflows, which often rely on manual and disjointed methods.
@@
 Using NLP, it efficiently explores this space, quickly identifying transformations and pragma configurations to achieve designs with good QoR.
-Our evaluation demonstrates that Sisyphus outperforms AutoDSE, NLP-DSE, and ScaleHLS in terms of QoR.
+Our evaluation demonstrates that Sisyphus outperforms AutoDSE, NLP-DSE, and ScaleHLS in terms of quality of result.
+Acknowledgments -   This work was supported by the NSF award #CCF-2211557.
+It is also supported by CDSC industrial partners and the AMD [^1] J.
+Cong has a financial interest in AMD.
+/HACC Program.